### v1

# PertEval: Unveiling Real Knowledge Capacity of LLMs with Knowledge-Invariant Perturbations

Abstract
Expert-designed close-ended benchmarks serve as vital tools in assessing the knowledge capacity of large language models (LLMs). Despite their widespread use, concerns have mounted regarding their reliability due to limited test scenarios and an unavoidable risk of data contamination. To rectify this, we present PertEval, a toolkit devised for in-depth probing of LLMs’ knowledge capacity through knowledge-invariant perturbations. These perturbations employ human-like restatement techniques to generate on-the-fly test samples from static benchmarks, meticulously retaining knowledge-critical content while altering irrelevant details. Our toolkit further includes a suite of transition analyses that compare performance on raw vs. perturbed test sets to precisely assess LLMs’ genuine knowledge capacity. Six state-of-the-art LLMs are re-evaluated using PertEval. Results reveal significantly inflated performance of the LLMs on raw benchmarks, including an absolute 21% overestimation for GPT-4. Additionally, through a nuanced response pattern analysis, we discover that PertEval retains LLMs’ uncertainty to specious knowledge, potentially being resolved through rote memorization and leading to inflated performance. We also find that the detailed transition analyses by PertEval could illuminate weaknesses in existing LLMs’ knowledge mastery and guide the development of refinement. Given these insights, we posit that PertEval can act as an essential tool that, when applied alongside any close-ended benchmark, unveils the true knowledge capacity of LLMs, marking a significant step toward more trustworthy LLM evaluation.
footnotetext: Work done during Li’s internship at Alibaba Group, under the guidance of Hu.footnotetext: Corresponding author: qiliuql@ustc.edu.cn.
1 Introduction
Large language models (LLMs) are developing rapidly, and have shown excellent basic capabilities such as reasoning [1, 2], planning [3] and world knowledge [4] in real-world tasks. With the large-scale deployment of LLMs in more and more scenarios, including those safety-critical ones [5], evaluating the real capability of LLMs becomes a necessary and significant task. Knowledge capacity, i.e., the ability to acquire, retrieve and utilize knowledge to solve professional problems, is one of the core capabilities of LLMs. Existing benchmarks for assessing knowledge capacity rely on standardized tests using static, expert-designed datasets [6, 7, 8, 9, 10, 11]. These test datasets consist of multiple-choice questions that include question descriptions, goals, options, and correct answers. They encapsulate valuable domain-specific knowledge that LLMs need to comprehend. The knowledge capacity of LLMs can be directly gauged by their performance on these test datasets.
So far, static data-based benchmarks have successfully evaluated the relative knowledge capacity of LLMs in various domains [12]. However, despite the high quality of expert-designed test data, static data-based benchmarks still face significant challenges in accurately quantifying the true knowledge capacity of LLMs due to their limited test scenarios and the unavoidable risk of data contamination. First, static data-based benchmarks depend on close-ended, single-formed questions for evaluation, which differs a lot to complicated real-world scenarios. For instance, instead of merely selecting an option, some users prefer asking LLMs to judge the correctness of options, while others may directly request the LLMs to generate an answer without providing an option list. Each possible prompt style may affect the performance of LLMs. Furthermore, Schaeffer et al. [13] have demonstrated that using nonlinear or discontinuous evaluation methods could lead to inaccurate evaluation results of LLMs’ capacity. Therefore, to genuinely evaluate the knowledge capacity of LLMs, it is necessary and feasible to test their ability to adjust to a range of complex scenarios that mirror real-world conditions. Second, static data-based benchmarks face an inherent risk of data contamination. These benchmarks publish test data online to ensure transparency and reproducibility in the evaluation process. However, this practice provides opportunities for LLM developers to maliciously fine-tune their models to "memorize" the test data, thereby undermining the benchmark’s integrity. This can result in an overestimation of the LLMs’ knowledge capacity and significantly compromise the reliability of the evaluation results. Efforts to prevent data contamination have primarily focused on detecting such contamination [14, 15] and generating new test data [16]. Nonetheless, high-quality test data contains valuable knowledge that has not been fully leveraged. We contend that it is feasible to utilize this knowledge effectively to assess the true knowledge capacity of LLMs while mitigating the risk of data contamination.
To achieve these goals, we propose PertEval, a knowledge-invariant perturbation-based evaluation toolkit for unveiling real knowledge capacity of LLMs in static data-based benchmarks. This idea stems from an analogy between LLM evaluation and humans’ educational assessment. Similar to LLM evaluation, close-ended questions are prevalent in human evaluation due to their versatility, cost effectiveness and precision of measurement [17], and also suffers from multiple challenges such as cheating [18] and a lack of variety [17]. Solutions include but limited to item personalization [19] and the design of distractive options [20, 21]. Inspired by this, PertEval includes human-like knowledge-invariant perturbations to transform test data to various complicated forms, and transition analysis methods to evaluate and trace the change of the performance of LLMs in different test scenarios. Knowledge-invariant perturbations consist of the rewriter LLM-based content-level perturbation for largely changing text features of questions to prevent data contamination, and a series of format-level perturbations for comprehensively covering possible test scenarios in real-world conditions. Transition analysis methods consist of not only evaluation metrics for measuring real knowledge capacity of LLMs, but also response pattern analysis for tracing the cause of the change of LLM’s performance, which probes in-depth weaknesses of the knowledge capacity of LLMs and provide insights for the refinement of LLMs.
In experiments, we re-evaluate six state-of-the-art LLMs using PertEval. Transition analysis results first reveal a significant overestimation of the knowledge capacity of LLMs in MMLU, with an absolute 21% overestimation for GPT-4 and 35% overestimation for Gemini-1.0-Pro. Response pattern analyses further unveil that the overestimation for most LLMs is caused by a significant increment of the ratio of extra incorrect choices in response patterns, with an increment up to 9.7% for GPT-4 and 27.7% for Gemini-1.0-Pro. These results illustrate the potential rote memorization of the LLMs on correct options. Detailed transition analyses in terms of overall performance stability and correct response consistency reveal various weaknesses in the knowledge capacity of existing LLMs, providing guidance for their refinement. Given these insights, we posit that PertEval serves as an essential tool for evaluating the real knowledge capacity of LLMs when applying to any existing close-ended benchmarks, marking a significant step towards more trustworthy LLM evaluation.
2 Related Work
Benchmarks for Knowledge Capacity Evaluation. Benchmarks for knowledge capacity evaluation could be classified into general knowledge benchmarks or professional knowledge benchmarks [12]. General knowledge benchmarks aim to evaluate the general knowledge capacity of LLMs in all kinds of domains, such as MMLU [6], C-Eval[22] and ARC[8]. On the other hand, professional knowledge benchmarks aim to deeply evaluate the professional capacity of LLMs in specific domains, such as MedMCQA [9] (Medicine), ScienceQA [23] (Science) and GSM-8K [11] (Mathematics). These benchmarks depend on professional multiple-choice questions to measure the capacity of LLMs.
Trustworthiness of Knowledge Capacity Evaluation. Research about the trustworthiness of knowledge capacity evaluation are mainly from the test data acquisition perspective and the response generation perspective. For test data acquisition, the data contamination risk have been proved widely exist in test data [24, 25, 26]. Recent works emphasize the detection of contamination [14, 15] and generating contamination-free test data [16, 27]. For response generation, it has been observed that multiple text-related biases could affect LLMs’ performance [28, 29, 30, 31], such as option ordering bias [28] and selection bias [29]. Debiasing methods like PriDe [14] serve as external plug-in for LLMs in multiple-choice question answering, which are inappropriate for evaluating real knowledge capacity.
Adversarial Text Attack for LLMs. Adversarial text attacks aim to mislead language models to generate wrong outputs. Important research directions of them in the era of LLM include but not limited to jailbreak attack and text classification attack. Traditional text attacks could be classified into character-level [32, 33], word-level [34] and sentence level [35]. Recently, LLM-based perturbations such as PromptAttack [36] have also been introduced. However, considering their cost and the comprehensiveness, existing text attack methods are insufficient for achieving our research goal.
3 Methodology
Table 1: Standards of knowledge invariance
Standard Name Standard Description
Semantic Information Invariance The perturbed question must have the same semantic information as the original question, which cannot change the name of entities, logic of statements and meaning of equations.
Reasoning Invariance A human test-taker’s reasoning process to obtain his/her response in the perturbed question should be consistent with that in the original question.
Answer Invariance The answer of a perturbed question should be semantically equivalent to the answer of the original question.
Statement Clarity The perturbed question should clearly present contexts, conditions and the target of the question.
3.1 Knowledge-invariant Perturbation
3.1.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
The content-level perturbation aim to largely change the statement of questions while keeping the original knowledge to prevent data contamination of orginal test data. The key challenge of knowledge-invariant paraphrasing is how to preserve the original knowledge while change the statement as possible it can. To this end, we first propose standards of knowledge invariance, as presented in Table 1 for the checking of knowledge invariance. Next, we propose a sentence-by-sentence paraphrasing algorithm using an LLM rewriter. The motivation is that the logic of a question text is the composition of the logic of the sentence sequences. Formally, let $q=(s_{1},s_{2},\ldots,s_{T})$ be the question text, where $s_{t},t=1,2,\ldots,T$ denotes the $t$-th sentence. Then the semantic of $s_{t}(t\geq 2)$ depends on its prerequisite sequence $(s_{1},\ldots,s_{t-1})$. To rewrite the whole question text, an LLM rewriter is instructed to rewrite sentence by sentence given each sentence and its original context. The paraphrasing algorithm, the prompt template for the rewriter LLM and an example of the perturbation are available at Appendix C.1.
3.1.2 Format-level Perturbation: Question Format Refactoring
Question format refactoring aim to evaluate the robustness of LLMs’ knowledge capacity when confronting complicated test conditions. To this end, we design a variety of format-level perturbation strategies that cover most possible format bias of LLMs in real-world scenarios.
Option Permutation (OptionPerm). OptionPerm reorders option contents while keeping the order of option IDs. It aims to evaluate the option ordering bias in the knowledge acquirement of LLMs. The default setting of OptionPerm is to reverse the order of option contents to completely destroy the local order of contents. When applying OptionPerm, the true answer of the question is also changed to be aligned with the perturbation result.
Option Format Refactoring (OptionForm). OptionForm changes the format of option IDs (e.g., append a right parenthesis to the end of IDs). It aims to evaluate the option format bias in the knowledge acquirement of LLMs. OptionForm might affect the performance of LLMs by changing the dependency beteween different tokens in options.
Option ID Shifting (OptionCaesar). OptionCaesar shifts the ASCII value of option IDs to change their character, which is similar to Caesar encryption. It aims to evaluate the selection bias in the knowledge acquirement of LLMs, whose existence has been empirically proved in some LLMs. By applying OptionCaesar, we replace common option IDs (e.g., A/B/C/D) by uncommon option IDs so that we can observe whether the value of IDs affect the performance of LLMs.
Question Type Changing (ChangeType). ChangeType changes a multiple choice question to a multiple judgement question. It aims to evaluate the question type bias of LLMs. Since the feasible solution space of a multiple choice question and the corresponding multiple judgement question is identical (given $N$ options, the size of the feasible solution space is $(2^{N}-1)$), an LLM that robustly acquire knowledge/skills in a question should be insensitive to ChangeType. Therefore, it is a significant perturbation for LLMs’ knowledge acquirement evaluation.
Question Position Swapping (SwapPos) SwapPos swaps the position of question texts with options. It aims to evaluate the global ordering bias of LLMs. For rational human test-takers, SwapPos does not change their performance on questions because merely move question text to after the options does not change semantic information and knowledge of the question. However, although token-level knowledge is also kept in SwapPos, this perturbation still might be a difficult challenge for LLMs, because it destroys the global ordering of input prompts, which might largely change the distribution of output of self-regression text generation models.
3.2 Knowledge Invariance Checking
LLM-based Knowledge Invariance Scoring. This method checks the knowledge invariance of perturbations from the perspective of LLM’s perception. Specifically, we first choose a superior LLM (e.g., gpt-4-turbo) as the referee. Then, give the test dataset $D_{dual}=\{(q_{i},q_{i}^{\prime})|i=1,2,\ldots,|D|\}$, we construct the scoring prompt based on the prompt template, standards of knowledge invariant perturbations (see Table 1) and the criteria of scores for knowledge invariance judgement (see Table 8 in Appendix C.1) for each pair of original and perturbed questions. After collecting the output score of the referee LLM from the scoring prompt, the average score is calculated as the final result.
Testing on Mastered Questions for LLMs. Different from the scoring method that is input-oriented, this method evaluates the output-oriented knowledge invariance. The motivation is that if a perturbation is knowledge-invariant, an LLM’s performance on questions whose knowledge it really mastered should be the same on the original and the perturbed version. The challenge is that mastered questions for LLMs are hard to find. To this end, we propose a substitute solution that utilizes questions that most LLMs can correctly answer to represent the set of mastered questions for an LLM. Then an LLM test-taker (gpt-4-turbo in experiments) is required to answer the original and perturbed versions of these questions to validate the knowledge invariance of the perturbation. In a word, the test-based knowledge invariance checking procedure can be summarized as follows:
  1. 1.
    Given a set of LLM test-takers and the evaluation dataset $D$, collect question samples that all LLMs answer correctly to construct the mastered question set $D_{simple}$.
  2. 2.
    Apply a perturbation to each of sample of $D_{simple}$ to obtain the perturbed dataset $D_{simple}^{\prime}$.
  3. 3.
    For each LLM, evaluate its performance on $D_{simple}^{\prime}$. If there exists at least one LLM whose performance on $D_{simple}^{\prime}$ has no significant difference with that on $D_{simple}$, then the perturbation is knowledge-invariant.
This method focuses more on the influence of perturbation on the text generation process of LLMs. Compared to knowledge invariance scoring operations, this methods is more suitable for evaluating the knowledge invariance of content-level perturbations.
3.3 Transition Analysis For Measuring Real Knowledge Capacity
The transition analysis aims to measure the real knowledge capacity of LLMs from their responses to benchmark datasets. In this part, we first propose the metric of real knowledge capacity. Next, we propose two aspects for detailed analyses of LLMs’ weaknesses in knowledge capacity.
Metric of Real Knowledge Capacity. To measure real knowledge capacity of LLMs, we propose the Consistent Accuracy (ACC@Consist) as the evaluation metric. The motivation is that if a question and its corresponding knowledge is truly mastered by an LLM, then the LLM’s performance should be consistent on any perturbed version of the question, including the original version and the most complicated version. To this end, let $M(\cdot)$ denotes the response function of an LLM. Let $x=(q_{x},y_{x})\in X$ denotes a test question, where $q_{x}$ denotes the question text and $y_{x}$ denotes the correct answer(s). Let $\sigma^{*}:X\to X$ denotes the most complicated composite knowledge-invariant perturbation. Let $D=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the test set. Then ACC@Consist is defined as:
[…]
Here $I(\cdot)$ denotes the indicator function.
Aspect 1: Overall Performance Stability. In terms of overall performance, we expect LLMs that have robustly acquired knowledge/skills required by benchmark datasets should have a stable performance when confronting knowledge-invariant perturbations. To this end, we propose Performance Drop Rate (PDR) to measure the over performance stability of LLMs given knowledge-invariant perturbations:
$\text{PDR}(M,D,\sigma)=\frac{1}{|D|}\sum_{x\in D}I[M(q_{\sigma(x)})=y_{\sigma(x)}]-I[M(q_{x})=y_{x}].$ (2)
Here $M$ denotes the LLM, $D$ and $D^{\prime}$ denotes the original and the perturbed dataset respectively. $PDR$ essentially equals to the discrepancy between the LLM’s accuracy on $D$ and that on $D^{\prime}$. If $PDR<0$, then the perturbation decreases the overall performance of LLMs. Then the LLM does not robustly acquire knowledge/skills. To further obtain reliable conclusions, we further make Wilcoxon signed-rank test for original and perturbed question sample pairs.
Table 2: Performance Transition Matrix
Number Correct after perturbation $\sigma$ Incorrect after perturbation $\sigma$
Correct before perturbation $\sigma$ $CC$ $IC$
Incorrect before perturbation $\sigma$ $IW$ $CW$
Aspect 2: Correct Response Consistency. To this end, we propose the Recall of Performance (ROP) metric. Its name originates from an analogy to the classical Recall score. To define ROP, we first propose the performance transition matrix, which is similar to the confusion matrix, as shown in Table 2. Here the first capital character of each variable (i.e., C/I) denotes “consistent” or “inconsistent”. The second capital character of each variable (i.e., C/N) denotes “correct” or “wrong”. For example, the variable $CC$ means the number of “consistent correct” before and after the perturbation. Next, similar to recall score, ROP is defined as the ratio of consistent correct in the number of all correct samples before perturbation, i.e.,
$\text{ROP}(M,D,\sigma)=\frac{CC}{CC+IC}.$ (3)
Tha range of ROP is in $[0,1]$. The higher the ROP, the better the correct response consistency given the perturbation. It should be noticed that, unlike PDR, there does not exist a “threshold” for ROP to measure the correct response consistency. Rather, the ROP serves as a score of correct response consistency, which can be utilized to compare the performance of different LLMs and strategies.
(a) Knowledge invariance score rated by GPT-4.
(b) Logarithm of Levenshtein distance.
Figure 1: LLM-based knowledge invariance scoring using systematic sampling (interval = 10).
Table 3: Results of testing on mastered questions - Performance Drop Rate (PDR) of overall performance stability testing on mastered questions using gpt-4-turbo.
Strategy C-Math W-History P-Psychology P-Medicine AVGmacro
KnInvPara $0$0.0115 0.0091 -0.0244 -0.0067
OptionPerm 0.0000 0.0000 0.0000 -0.0244 -0.0061
OptionForm 0.0000 0.0000 0.0000 0.0000 0.0000
OptionCaesar 0.0000 0.0000 0.0091 0.0000 0.0023
ChangeType 0.0000 0.0000 0.0000 0.0000 0.0000
SwapPos $0$0.1149 -0.0636 -0.0488 -0.0568
4 Experiment
Datasets. In the experiment, to comprehensively cover evaluation domains and subjects while keeping the evaluation cost, we select test data in College Mathematics (C-Math), World History (W-History), Professional Psychology (P-Psychology) and Professional Medicine (P-Medicine) from the Massive Multitask Language Understanding (MMLU) benchmark, with each representing a supercategory in STEM, Humanities, Social Sciences and Other respectively. Statistics of selected datasets are presented in Table 9 in Appendix D.
Large Language Models. Based on existing LLM leaderboards like OpenCampass [^1] https://rank.opencompass.org.cn/ and considering both the popularity and timeliness of LLMs, we select six representative LLMs for evaluation. These LLMs include close-sourced (gpt-4-turbo, gpt-3.5-turbo, gemini-1.0-pro, glm-3-turbo) and open-sourced (mistral-7b-instruct-v0.2, llama-3-8b-instruct).
4.1 Checking Knowledge Invariance for Perturbations
Knowledge invariance checking results are presented in Figure 1 and Table 3. For LLM-based knowledge invariance scoring (Figure 1), we compare the score of proposed perturbations with the baseline PromptAttack[36] and find the excellent average score ($>4.5$) for each proposed perturbation. We also calculate the Levenshtein distance (edit distance) of perturbed and original questions for each perturbation, and we find that the traditional edit distance is not suitable for measuring knowledge invariance because of a lack of the correlation between edit distance and knowledge invariance score. For testing on mastered questions (Table 3), we find that most perturbations obtain an average PDR closed to zero, which indicates their consistent performance on the perturbed and the original data. The PDR of SwapPos on World History and Professional Psychology is less than -0.05 although it does not change any knowledge-relevant information based on our design. A possible explanation is that SwapPos changes the global ordering of question prompts, thus affects the text generation of these self-regression-based LLMs. More detailed result analyses are available at Appendix D.1
4.2 Real Knowledge Capacity Evaluation (Transition Analysis Part 1)
Figure 2: Real knowledge capacities measured by ACC@Consist with composite knowledge-invariant perturbation. Metrics are averaged on datasets. ACC@Original denotes accuracy from the original data. ACC@Perturb denotes accuracy from the perturbed data. ACC@Consist denotes ratio of consistent correct responses in both the original and the perturbed data.
In PertEval, each knowledge-invariant perturbation strategy focuses on a specific question transformation. To construct the most challenging test scenario for LLMs and obtain reliable and conprehensive evaluation results, we compose all knowledge-invariant perturbations together to obtain the composite perturbation to quantify LLMs’ real knowledge capacity. Then the real knowledge capacity of LLMs is measured by ACC@Consist. Evaluation results are shown in Figure 2. Based on these results, we obtain several findings in terms of LLMs’ knowledge capacity evaluation:
1. The knowledge capacity of LLMs is significantly overvalued in static dataset-based benchmarks. Comparing the ACC@Original with ACC@Perturb and ACC@Consist for each LLM, we observe a significant drop of LLMs’ performance on perturbed dataset. Especially for gpt-4-turbo, gpt-3.5-turbo and gemini-1.0-pro, the change exceeds 20%. Even the most powerful selected LLM, gpt-4-turbo, gains only ACC@Consist as 0.5551. These results indicate a large gap between evaluated knowledge capacity of LLMs in static data-based benchmarks and real knowledge capacity of LLMs in complicated scenarios.
2. Every selected LLM has its own consistently mastered knowledge. This result is deduced from a comparison between LLMs’ performance and pure guess. Given $k$ options and a single answer, the expected ACC@Consist is $1/k^{2}$. Therefore, the expected of ACC@Consist for pure guess is $0.06251/4^{2}=$ 0.0625 in the experiment ($k=4$)222see Appendix D.2 for detailed demonstration.. We observe from Figure 2 that ACC@Consist values of all selected LLMs are significantly larger than 0.0625, which means that each selected LLM indeed has its own consistently mastered knowledge, although most of them master less than a half (ACC@Consist < 0.5) of the total knowledge.
4.3 Response Pattern Analysis (Transiton Analysis Part 2)
(a) Original data
(b) Perturbed data with composite perturbation
Figure 3: Response patterns of gpt-4-turbo in all datasets.
We conduct response pattern analysis to explore how perturbations affect the performance of LLMs, and we discover that knowledge-invariant perturbations affect LLMs by increasing the ratio of extra multiple choices. As presented in Figure 3, the main reason why the performance of gpt-4-turbo drops dramatically is that they select extra incorrect options more frequently in the perturbed data (gpt-4-turbo: $12\%\to 21.7\%$). These LLMs tend to select extra incorrect options besides selecting correct options. A possible explanation of this phenomenon is that the LLMs memorize correct options, while lack the ability to recognize and filter out incorrect options, which means that these LLMs do not really master knowledge of the questions. Therefore, besides the ability to distinguish correct answers, we should also emphasize enhancing large language models’ ability to recognize and filter out incorrect answers when facing options in real-world problems. More detailed response pattern analysis results are available at Appendix D.3.
4.4 Overall Performance Stability (Transition Analysis Part 3)
This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of overall performance. Evaluation results of overall performance stability are shown in Table 4. Here each value in the table’ cell is the macro PDR, i.e., the average value of PDRs separately calculated on each dataset. The Wicoxon signed-rank test is applied to the micro PDR, i.e., the overall PDR calculated on the all-in-one dataset. Specifically, let $s=I(M(q_{x})=y_{x})\in\{0,1\}$ and $s^{\prime}=I(M(q_{\sigma(x)})=y_{\sigma(x)})\in\{0,1\}$ respectively denotes model $M$’s score on the original and the perturbed questions, the alternative hypothesis of the Wilcoxon signed-rank test is $H_{a}:s>s^{\prime}$, or, $H_{a}:s^{\prime}-s<0$. From Table 4, we can both analyze for each column or for each row. For columns (strategy-aspect), we first observe that all selected LLMs shows negative macro PDRs given the content-level perturbation, the knowledge-invariant paraphrasing (KnInvPara). In addition, for close-sourced LLMs, the micro PDR is significantly negative, which means that the these LLMs significantly lack the content-level knowledge acquirements robustness in these datasets. Another observation is that all the LLMs are highly sensitive to the SwapPos perturbation. We suppose the reason is that, unlike other format-level perturbations that mostly focus on local format, SwapPos changes the global format of questions, which highly destroys self regression-based LLMs’ text generation process, even though this perturbation is entirely knowledge-invariant.
Table 4: Macro PDR $\uparrow$ and hypothesis test results of Micro PDR of LLMs w.r.t. perturbation.
[…]
gpt-4-turbo -$0∗$0208∗∗ -0.0136 -0.0294∗∗ -0.0210 -0.1117∗∗ -0.0468
gpt-3.5-turbo -0.0275∗∗ -0.0042 -0.1767∗∗ -0.0396∗∗ -0.1736∗∗ -0.1943∗∗ -0.1047
gemini-1.0-pro -0.0558∗∗ +0.0121 +0.0125 +0.0030 -0.1310∗∗ -0.1532∗∗ -0.0521
glm-3-turbo -$0∗$0190 -0.1397∗∗ -0.0118 +0.0522 -0.2142∗∗ -0.0667
mistral-7b-instruct-v0.2 -0.0264 -$0$0.2789∗∗ +0.0793 -0.0844∗∗ -0.1275∗∗ -0.0763
llama-3-8b-instruct -0.0336∗∗ -0.0091 -0.0939∗∗ -0.0368∗∗ -0.2920∗∗ -0.1814∗∗ -0.1074
AVG -0.0456 -0.0087 -0.0994 0.0003 -0.0596 -0.1335
  • ∗∗: The Micro PDR is significantly negative in the Wilcoxon signed-rank test ($\alpha=0.01$).
Table 5: Macro Recall of Performance (ROP) $\uparrow$ of LLMs w.r.t. perturbation strategies.
[…]
gpt-4-turbo 0.8798 0.9063 0.9601 0.9349 0.9221 0.7995 0.8949
gpt-3.5-turbo 0.8230 0.7728 0.5602 0.7801 0.5194 0.5610 0.6631
gemini-1.0-pro 0.7968 0.7995 0.9553 0.9226 0.6788 0.6362 0.7982
glm-3-turbo 0.7164 0.6756 0.5943 0.7884 0.6816 0.4026 0.6288
mistral-7b-instruct-v0.2 0.6543 0.6202 0.1339 0.7679 0.5128 0.4060 0.5158
llama-3-8b-instruct 0.8055 0.7433 0.7541 0.8102 0.3971 0.5209 0.6676
AVG 0.7522 0.7624 0.6673 0.8323 0.6858 0.6009
4.5 Correct Response Consistency (Transition Analysis Part 4)
This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of the consistency of correct responses. Experimental results of ROP are presented in Table 5. First, in terms of LLMs, the correct response consistency of different LLMs vary a lot. Among them, gpt-4-turbo performs best with the average macro ROP as 0.8949, which is a big gap compared to other LLMs. We also observe that gemini-1.0-pro performs well especially on the OptionForm and OptionCaesar strategies (Macro PDR > 0.9), which can detect the selection bias of LLMs. Without prerequisite information, we suppose that special effort have been made in the training of gemini-1.0-pro to overcome selection bias. For other LLMs, their correct response consistencies are vulnerable to almost all knowledge-invariant perturbations, which means that their correct answer in the original datasets are less reliable. On the other hand, the effect of different strategies on correct response consistency also vary a lot. Similar to results in overall performance stability, SwapPos is the most powerful strategy that influences LLMs’ correct response consistency, which exposes the inherent flaw of LLMs in defending global order-level perturbation.
5 Conclusion
In this paper, we proposed PertEval, an evaluation toolkit for measuring large language models’ real knowledge capacity using human-like knowledge-invariant perturbations and transition analysis methods. Based on the well-defined knowledge perturbation strategies and metrics in transition analysis, we revealed the overestimated knowledge capacity of six state-of-the-art LLMs on the MMLU benchmark. Response pattern analyses discovered that PertEval retains LLMs’ uncertainty to specious knowledge and revealed their rote memorization to correct options. Further transition analysis by PertEval illuminated the vulnerability of different LLMs to different perturbations and could guid the development of refinement. Finally, we discussed ways to enhance the knowledge capacity of LLMs using PertEval.
Appendix
Appendix A Related Work
A.1 LLMs’ Knowledge Capacity Evaluation Benchmarks
LLMs’ knowledge capacity evaluation plays a fundamental yet significant role in LLM evaluation tasks. The knowledge capacity evaluation aims to quantify LLMs’ ability to master and utilize professional knowledge and skills to solve real-world problems. Therefore, existing benchmarks usually depend on expert-designed test datasets, especially multiple choice questions, to evaluate LLMs’ knowledge capacity.
In terms of general knowledge benchmarks, MMLU [6] proposed to comprehensively evaluate LLMs’ professional knowledge on 57 subjects using the classical multiple choice question-based test. Similarly, C-Eval [22] focuses on evaluating LLMs’ advanced knowledge and reasoning abilities in Chinese context. AGIEval [7] collects human-centric datasets from real-world examinations such as SAT, GRE and GMAT to evaluate LLMs’ knowledge capacity in human-centric tests. Differently, ARC [8] measures LLMs’ knowledge capacity from knowledge styles rather than knowledge domains, including but not limited to “Definition”, “Basic Facts & Properties”, “Algebratic”, etc. In a word, general knowledge benchmarks usually consist of a knowledge taxonomy and a series of corresponding labelled datasets, each of which focusing on evaluating a specific aspect or domain of knowledge capacity for LLMs.
In terms of specific and professional knowledge benchmarks, expert-designed datasets are also the top priority. For clinical knowledge, MedMCQA [9], and PubMedQA [37] utilize professional multiple choice questions to assess LLMs’ clinical knowledge capacity and disease diagnosis ability. For social science, Ziems et al. [38] explored evaluating LLMs’ knowledge capacity in computational social science using real-world classification tasks. Nay et al. [10] explored LLMs’ knowledge capacity in tax law with a case study using automatically generated multiple choice questions. For science and engineering subjects, Arora et al. [39] proposed JEEBench, a challenging scientific problem solving benchmark consisting of mathematics, physics and chemistry problems. Other representative benchmarks such as GSM-8K [11] and ScienceQA [23], also utilize expert-designed professional test data to assess LLMs knowledge capacity from various perspectives.
A.2 Trustworthiness of Knowledge Capacity Evaluation
Large language models are self-regressive text generation models trained on large-scale online-accessible textual data. Therefore, from both the test data acquisition perspective and the response generation perspective, there exist various risks in the trustworthiness of LLMs’ knowledge capacity evaluation. These risks highly threaten the correctness and reliability of the knowledge capacity evaluation results.
From the test data acquisition perspective, the test data contamination problem has been proved widely exist in evaluation benchmarks [24, 25, 26]. Test data contamination can lead to overvalued knowledge capacity, and should be avoided at all costs. To this end, much effort have been devoted into distinguishing contaminated data or preventing data contamination with new test data. Magar and Schwartz [40] studied the problem for BERT models, and proposed a principled method to analyze the exploitation and memorization of contaminated data. Jacovi et al. [26] suggested using pre-evaluation intervention such as encrypting test data and avoiding uploading textual data to the internet to prevent data contamination. On the other hand, Bai et al. [27] proposed to utilize an examiner LLM to generate test questions and evaluate LLMs’ knowledge capacity to avoid data contamination. Recently, Golchin and Surdeanu [15] and Oren et al. [14] respectively proposed multi-level and example order-inspired methods for identifying test data contamination. In a word, existing works usually prevent data contamination by avoiding testing on contaminated data or generating brand-new test data. However, how to avoid data contamination while sufficiently utilize valuable knowledge information of existing test data for evaluation is still underexplored.
From the response generation perspective, LLMs’ biases towards specific prompt features in knowledge capacity evaluation have been widely observed [28, 29, 30, 31] which can highly influence their performance. Pezeshkpour and Hruschka [28] discovered that LLMs’ output sensitivity to specific ordering of options. For instance, some LLMs prefer selecting the first option in the option list, which proved the ordering bias in multiple choice question-based evaluation. Li et al. [30] further proposed methods to quantify consistence and confidence of LLMs’ accuracy on test datasets. Khatun and Brown [31] explored the usefulness of multiple choice question-based dataset for evaluating open-source LLMs, and found that many small-scale open-source LLMs fail to properly understand and select an answer from given choices. Recently, Zheng et al. [29] detected the selection bias towards options in multiple choice question answering for LLMs, and proposed PriDe, a label-free and inference time-free debiasing method. However, as an external plug-in for LLMs, this method is inappropriate for application to measuring the knowledge capacity of large language models themselves. In summary, existing research usually emphasize empirically detecting biases in the response generation procedure and proposing debiasing methods. Howver, how to avoid such bias to evaluate LLMs’ real knowledge capacity remains a research question.
Appendix B Discussion about PertEval
The ultimate goal of PertEval is not only to unveil the real knowledge capacity of LLMs, but also to guide the refinement of LLMs in enhancing knowledge capacity. To this end, we discuss potential methods and future directions for enhancing LLMs’ knowledge capacity with PertEval:
Supervised fine-tuning with perturbed data. PertEval can be utilized to generate various forms of multiple-choice question data. For knowledge that requires memorization such as history and laws, new data generated by PertEval can be used to enhance the ability of LLMs to adjust to various complicated test scenarios. Indeed, we conduct a primary experiment using llama-3-8b-instruct (see Appendix D.4) and reveal the feasibility of this solution.
Extending PertEval to open-ended test data. The current PertEval adapts only to multiple-choice questions considering their precision of measurement and cost effectiveness. To further evaluate the knowledge capacity of LLMs in all kinds of domains, a future direction is to extend PertEval to open-ended data types such as open question answering. Potential challenges in this direction include but not limited to how to define the correctness of output answer given perturbations and how to ensure knowledge invariance in complicated questions such as solving math application questions.
Appendix C Details of Methodology
C.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
The algorithm of knowledge-invariant paraphrasing using LLM rewriter is presented in Algorithm 1. The prompt template for the LLM rewriter is presented in Figure 4. In implementation, an expected similarity score is provided to the LLM rewriter to control the degree of the change of question text. In experiments, the expected similarity score is fixed to 0.6. An example of knowledge-invariant paraphrasing in MMLU college mathematics is shown in Table 6.
Algorithm 1 Knowledge-invariant Paraphrasing Using LLM Rewriter
1:Input: $q=(s_{1},s_{2},\ldots,s_{T})$, the question text; $M$, the LLM rewriter
2:Output: $q^{\prime}=(s_{1}^{\prime},s_{2}^{\prime},\ldots,s_{T}^{\prime})$, the perturbation output
3:procedure knowledgeInvariantParaphrase($q,M$)
4:     Initialize rewriting prompt template $p_{template}(\cdot)$
5:     $q^{\prime}\leftarrow\emptyset$
6:     for $t\in 1,2,\ldots,T$ do
7:         $p_{current}\leftarrow p_{template}(s_{t},(s_{1},\ldots,s_{t-1}))$triangleright Generate the rewriting prompt
8:         $r_{current}\leftarrow M.request(p_{current})$triangleright Obtain the LLM rewriter’s output
9:         $s_{t}^{\prime}\leftarrow sentenceFilter(r_{current})$
10:         $q^{\prime}\leftarrow q^{\prime}\oplus s_{t}^{\prime}$triangleright Append the perturbed sentence to the end of the question
11:     end for
12:     return $q^{\prime}$
13:end procedure
Here is a sentence in a multiple choice question. Please rewrite the sentence given its context and the expected similarity score. Here are necessary requirements: [Requirements Start] 1. Be consistent with its context. 2. The rewrited sentence should keep the semantic of the original sentence. 3. If the sentence contains blanks/underlines to be filled, these blanks/underlines should be kept after paraphrasing. 4. You can utilize various rewriting skills (e.g., add/replace/delete words, paraphrase) to make it looks different from the original. [Requirements End]
[Meaning of Expected Similarity Score Start] For the expected similarity score (0.0 - 1.0), 1.0 denotes that the rewrited is exactly the same as the original; 0.8 denotes that the the there exist word-level differences between the rewrited and the original; 0.6 denotes that there exist not only word-level, but lots of sentence structure-level differences between the rewrited and the original; 0.4 denotes that you are allowed to entirely paraphrase the sentence by your own; 0.2 denotes that you are allowed to add misleading statements to the current sentence. [Meaning of Expected Similarity Score End]
You should only output the rewrited sentence without any extra content. Expected similarity score: {similarity_score} Context: {context} Sentence: {sentence} Your output:
Figure 4: Prompt template for the rewriter LLM.
Table 6: An example of knowledge-invariant paraphrasing of a test question. Texts surrounded by angular brackets are invisible in question prompts input to the LLM test-taker.
Original Knowledge-invariant Paraphrasing
<# Context & Condition> Let $T:R^{2}\to R^{2}$ be the linear transformation that maps the point (1, 2) to (2, 3) and the point (-1, 2) to (2, -3).
<# Goal> Then $T$ maps the point (2, 1) to
<# Context & Condition> Let $T$ be the linear transformation from $R^{2}$ to $R^{2}$ such that $T$ maps (1, 2) to (2, 3) and (-1, 2) to (2, -3).
<# Goal> Then, the linear transformation $T$ will map the point (2, 1) to
C.2 Format-level Perturbation: Question Format Refactoring
Examples of question format refactoring are shown in Table 7.
Table 7: Examples of format-level knowledge-invariant perturbations. Texts surrounded by angular brackets are invisible in question prompts input to the LLM test-taker.
Perturbation Original case Perturbed case
OptionPerm <# Options>
A $x=1$; B $x=2$; C $x=3$; D $x=4$
<# Options>
A $x=4$; B $x=3$; C $x=2$; D $x=1$
OptionForm <# Options>
A $x=1$; B $x=2$; C $x=3$; D $x=4$
<# Options>
A) $x=1$; B) $x=2$; C) $x=3$; D) $x=4$
OptionCaesar <# Options>
A $x=1$; B $x=2$; C $x=3$; D $x=4$
<# Options>
U $x=1$; V $x=2$; W $x=3$; X $x=4$
ChangeType <# Prompt> Please select correct option(s) given the following question: <# Prompt> Please judge whether each of the options is correct given the following question:
SwapPos <# Prompt> Please select correct option(s) given the following question:
<# Question> The solution of the equation $32x+1=$ 3 is
<# Options> A $x=1$; B $x=2$; C $x=3$; D $x=4$
<# Prompt> Please select correct option(s) given the following question:
<# Options> A $x=1$; B $x=2$; C $x=3$; D $x=4$
<# Question> The solution of the equation $32x+1=$ 3 is
Table 8: The criteria of scores for knowledge invariance judgement.
Score Criteria
5.0 The perturbation perfectly satisfies all the standards and is entirely acceptable.
4.0 There are only minor flaws in the perturbed question. As a whole, the perturbed question is acceptable.
3.0 Only some parts of the perturbation is acceptable. As a whole, the perturbed question is less acceptable.
2.0 There are major flaws in the perturbed question that makes it unacceptable.
1.0 There are fatal flaws in the perturbed question that makes it entirely unacceptable.
Your task is to grade the knowledge invariance degree of a perturbed multiple choice question against the original question. You clearly know that if a perturbed question is knowledge-invariant, the question has to satisfy the following requirements: [Perturbation Requirements Start] 1. Semantic Information Invariance. The perturbed question must have the same semantic information as the original question, which cannot change the name of entities, logic of statements and meaning of equations. 2. Reasoning Invariance. A human test-taker’s reasoning process to obtain his/her response in the perturbed question should be consistent with that in the original question. 3. Answer Invariance. The answer of a perturbed question should be semantically equivalent to the answer of the original question. 4. Statement Clarity. The perturbed question should clearly present contexts, conditions and the target of the question without ambiguous statement. [Perturbation Requirements End]
The grading score is from 1 to 5. Grading criteria are given in the following: [Grading Criteria Start] 1.0 - There are fatal flaws in the perturbed question that makes it entirely unacceptable. 2.0 - There are major flaws in the perturbed question that makes it unacceptable. 3.0 - Only some parts of the perturbation is acceptable. As a whole, the perturbed question is less acceptable. 4.0 - There are only minor flaws in the perturbed question. As a whole, the perturbed question is acceptable. 5.0 - The perturbation perfectly satisfies all the requirements and is entirely acceptable. [Grading Criteria End]
[Original Question Start]: {original_question} [Original Question End]
[Perturbed Question Start]: {perturbed_question} [Perturbed Question End]
Figure 5: Prompt template for the LLM-based knowledge invariance scoring (Part 1/2).
You should grade the perturbation following these steps: 1. Recall the perturbation requirements and grading criteria, and read the original and the perturbed questions in detail. 2. For each of perturbation requirements, carefully judge its satisfaction degree of the perturbed question. 3. Based on step 1 and step 2, give a total grading score for the perturbed question. 4. Analyze strengths and weakness of the perturbed question from the view of perturbation requirements based on step 1,2,3. Think carefully for a while, then propose your conclusion. Your output template is given as follows:
[Template Start] { "score": <numeric score from 1 to 5>, "strength": <"xxx", strengths of the perturbation>, "weakness": <"xxx", weaknesses of the perturbation> } [Template End]
Your conclusion:
Figure 6: Prompt template for the LLM-based knowledge invariance scoring (Part 2/2).
Appendix D Details of Experiments
Statistics of datasets are presented in Table 9.
Table 9: Dataset Statistics
Name College Mathematics High School World History Professional Psychology Professional Medicine
Supercategory STEM Humanities Social Science Others
Concepts Differential equations, Ottoman empire, Diagnosis, Diagnosis,
real analysis, economic imperialism, biology and behavior, pharmacotherapy,
combinatorics… World War I… lifespan development, … disease prevention…
# Questions 100 237 612 272
# Tokens per Q 46.00$\pm$25.37 290.25$\pm$124.13 28.02$\pm$19.53 144.66$\pm$65.41
# Tokens per P 129.52$\pm$29.19 392.27$\pm$127.62 125.11$\pm$31.63 233.58$\pm$66.60
Figure 7: Accuracy score of LLMs in original datasets. The gpt-4-turbo@Mastered denotes the accuracy of gpt-4-turbo on the mastered question set for checking knowledge invariance, which is defined in Section 3.2.
D.1 Checking the Knowledge Invariance of Perturbation Strategies
In this part, we utilize the proposed knowledge invariance checking methods to measure the validity of perturbations in terms of knowledge invariance. The prompt template for the LLM-based knowledge invariance scoring is presented in Figure 5 and 6. Results of LLM-based knowledge invariance scoring are shown in Figure 1. Results of overall performance stability testing on mastered questions are presented in Figure 3. We obtain several conclusions from these experimental results, as given in the following:
C1. Proposed perturbations are knowledge-invariant in most cases. From the right part of Figure 1, we observe that knowledge invariance scores of our proposed perturbations exceeds 4.5 and exceeds the baseline in most cases. For the content-level KnInvPara, since paraphrasing question confronts inevitable semantic changing, its knowledge invariance score is not as high as that of format-level perturbations. Indeed, we can observe from Table 3 that PDRs of KnInvPara are always closed to zero, which demonstrates its potential knowledge invariance.
C2. Traditional string-based text distance metric is insufficient for measuring knowledge invariance. In the left part of Figure 1, we show the average logarithm of Leveinshtein distance (edit distance) for each perturbation. Comparing the Levenshtein distances of PromptAttack, global text ordering-level format perturbation (ChangeType, SwapPos) and KnInvPara, we find that although PromptAttack has the lowest Leveinshtein distance in most datasets, its knowledge invariance scores are also relatively low. On the other hand, PromptAttack have the highest Leveinshtein distance in W-History, but also has the lowest knowledge invariance score in the four datasets. Therefore, As a result, there lacks an obvious correlation between Leveinshtein distance and knowledge invariance score. Essentially, this result indicates that knowledge relevant features of texts could be be disentangled from string-level features to some extent. How to efficiently and effectively measure knowledge invariance or quantify knowledge relevant features of texts remains a research problem.
C3. LLMs are vulnerable to global text order-level knowledge-invariant perturbations even in mastered questions. We observe from Table 3 that SwapPos can decrease the performance of gpt-4-turbo even in mastered questions although we have demonstrated the knowledge invariance of the perturbation in Figure 1, Comparing PDRs on different datasets and statistics of different datasets, we observe that it is highly correlated with the number of tokens per question.
D.2 Probing LLMs’ Consistent Knowledge Capacity
Proposition D.1.
In multiple choice questions, given $k$ options and one single correct answer for each question, the expected value of ACC@Consist is $1/k^{2}$ for pure guessing.
Proof.
[…]
$E[\text{ACC@Consist}]=E[X/n]=\frac{1}{n}E[X]=1/k^{2}$ (4)
Then the proof is completed. ∎
D.3 Response Pattern Analysis for LLMs in PertEval
In this part, we analyze response pattern for LLMs in the original and perturbed data, which can provide us a deep insight into the influence of knowledge-invariant perturbations on LLMs’ question answering procedure. Specifically, response patterns consist of correct choice, invalid choice, extra multiple choice, wrong single choice and wrong multiple choice. Their definition are listed in the following:
  • Correct choice: The LLM selects and only selects all correct options. The ratio of correct choices is equvalent to micro accuracy.
  • Invalid choice: The LLM selects none of options. This could be caused by the LLM’s wrong reasoning process or invalid output format.
  • Extra multiple choice: The LLM selects not only all correct options, but also extra incorrect options. This means that the LLM fails to filter out incorrect options from the option list.
  • Wrong single choice: The LLM selects only a single option, and the option is wrong. This means that the LLM fails to recognize the correct option in the option list.
  • Wrong multiple choice: The LLM selects multiple options which do not cover all correct options.
Given the response pattern taxonomy, we counted the ratio of response patterns in original and perturbed data for all selected LLMs, as shown in Figure 3, 8, 9, 10, 11 and 12.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 8: Response patterns of gpt-3.5-turbo in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 9: Response patterns of glm-3-turbo in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 10: Response patterns of gemini-1.0-pro in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 11: Response patterns of llama-3-8b-instruct in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 12: Response patterns of mistral-7b-instruct-v0.2 in all datasets.
D.4 How to Enhance LLMs’ Knowledge Capacity Using PertEval?
The goal of PertEval is not only to evaluate large language models’ consistent knowledge capacity, but to help LLMs enhance their knowledge capacity in real-world applications. To this end, we explore utilizing perturbed data generated by PertEval to fine-tune the open-sourced LLM, LLaMA-8B-Instruct, as presented in Table 10. We obtain several insights from the experimental results.
For format-level perturbations, we recommend fine-tune LLMs on a small set of perturbed data to enhance their consistent knowledge capacity in all knowledge domains. This conclusion originates from the observed stimulation phenomenon in fine-tuning. That is, only fine-tuning the model with a subset of perturbed data can significantly improve its overall performance stability in all perturbed data. This result indicates the ability of LLMs to “acquire” format-level perturbations and perform consistently on perturbed data in all kinds of knowledge domains.
For content-level perturbations, we recommend training knowledge expert models using both the original and the perturbed data for specific knowledge domains. This conclusion orignates from the lack of transferability of content-level fine-tuning for datasets. This observation is reasonable because the knowledge and statement styles of different domain vary a lot. Considering the numerous number of knowledge domains and the issue of cost and efficiency, we suppose training knowledge expert models for specific knowledge domains is a good choice for striking a balance between knowledge capacity and training efficiency.
Table 10: PDR of LLaMA-8B-Instruct before and after finetuning on perturbed datasets. Underlined datasets are used for finetuning. CT denotes ChangeType. KP denotes KnInvPara. SP denotes SwapPos. For example, F(CT) means finetuning LLaMA-8B-Instruct on world history and professional medicine datasets perturbed by ChangeType.
[…]
ChangeType Original -0.2300 -0.3418 -0.2467 -0.3493 -0.2920 -0.1998
F(CT) -0.0700 +0.0759 +0.0196 +0.0257 +0.0128 -0.0868
F(CT+KP) -0.0500 +0.0422 +0.0082 +0.0074 +0.0020 +0.0019
SwapPos Original -$0$0.2110 -0.1944 -0.2500 -0.1814 -0.2867
F(SP) +0.0100 -0.0675 -0.1095 -0.0882 -0.0638 +0.0246
F(SP+KP) -$0$0.1350 -0.1029 -0.1176 -0.0964 -0.1065
KnInvPara Original +0.0200 -0.0802 -0.0163 -0.0478 -0.0311 -0.0328
F(KP) -$0$0.0253 -0.0212 $0$0.0216 -0.0188
F(CT+KP) -$0$0.0549 -0.0343 -0.0368 -0.0415 -0.0393
F(SP+KP) -$0$0.0675 -0.0212 -0.0184 -0.0343 -0.0303

### v2

@@
 Abstract
-Expert-designed close-ended benchmarks serve as vital tools in assessing the knowledge capacity of large language models (LLMs).
+Expert-designed close-ended benchmarks are indispensable in assessing the knowledge capacity of large language models (LLMs).
 Despite their widespread use, concerns have mounted regarding their reliability due to limited test scenarios and an unavoidable risk of data contamination.
@@
 These perturbations employ human-like restatement techniques to generate on-the-fly test samples from static benchmarks, meticulously retaining knowledge-critical content while altering irrelevant details.
-Our toolkit further includes a suite of transition analyses that compare performance on raw vs.
+Our toolkit further includes a suite of response consistency analyses that compare performance on raw vs.
 perturbed test sets to precisely assess LLMs’ genuine knowledge capacity.
-Six state-of-the-art LLMs are re-evaluated using PertEval.
-Results reveal significantly inflated performance of the LLMs on raw benchmarks, including an absolute 21% overestimation for GPT-4.
-Additionally, through a nuanced response pattern analysis, we discover that PertEval retains LLMs’ uncertainty to specious knowledge, potentially being resolved through rote memorization and leading to inflated performance.
-We also find that the detailed transition analyses by PertEval could illuminate weaknesses in existing LLMs’ knowledge mastery and guide the development of refinement.
-Given these insights, we posit that PertEval can act as an essential tool that, when applied alongside any close-ended benchmark, unveils the true knowledge capacity of LLMs, marking a significant step toward more trustworthy LLM evaluation.
-footnotetext: Work done during Li’s internship at Alibaba Group, under the guidance of Hu.footnotetext: Corresponding author: qiliuql@ustc.edu.cn.
+Six representative LLMs are re-evaluated using PertEval.
+Results reveal significantly inflated performance of the LLMs on raw benchmarks, including an absolute 25.8% overestimation for GPT-4.
+Additionally, through a nuanced response pattern analysis, we discover that PertEval retains LLMs’ uncertainty to specious knowledge, and reveals their potential rote memorization to correct options which leads to overestimated performance.
+We also find that the detailed response consistency analyses by PertEval could illuminate various weaknesses in existing LLMs’ knowledge mastery and guide the development of refinement.
+Our findings provide insights for advancing more robust and genuinely knowledgeable LLMs.
+Our code is available at https://github.com/aigc-apps/PertEval.
+footnotetext: Work done during Li’s internship at Alibaba Cloud Computing, under the guidance of Hu.footnotetext: Qi Liu is the corresponding author.
 1 Introduction
-Large language models (LLMs) are developing rapidly, and have shown excellent basic capabilities such as reasoning [1, 2], planning [3] and world knowledge [4] in real-world tasks.
-With the large-scale deployment of LLMs in more and more scenarios, including those safety-critical ones [5], evaluating the real capability of LLMs becomes a necessary and significant task.
-Knowledge capacity, i.e., the ability to acquire, retrieve and utilize knowledge to solve professional problems, is one of the core capabilities of LLMs.
-Existing benchmarks for assessing knowledge capacity rely on standardized tests using static, expert-designed datasets [6, 7, 8, 9, 10, 11].
-These test datasets consist of multiple-choice questions that include question descriptions, goals, options, and correct answers.
+Figure 1: An overview of the PertEval evaluation toolkit.
+PertEval uses content-level and format-level perturbations to generate perturbed dataset $D^{\prime}$ from existing close-ended benchmark dataset $D$.
+Next, it evaluates the knowledge capacity of LLMs via response consistency analysis.
+PertEval also demonstrates in-depth the performance feature of LLMs via response pattern analysis.
+Large language models (LLMs) are developing rapidly, and have shown excellent basic capabilities, such as reasoning [1, 2], planning [3] and world knowledge [4], in real-world tasks.
+Given the widespread deployment of LLMs across an increasing number of scenarios, including those that are safety-critical [5], evaluating the real capability of LLMs becomes a necessary and significant task.
+Knowledge capacity, i.e., the ability to retrieve and utilize acquired knowledge to solve professional problems, is one of the core capabilities of LLMs [4].
+Existing knowledge capacity evaluation largely rely on standardized tests using close-ended benchmarks [6, 7, 8, 9, 10, 11].
+These benchmarks consist of multiple-choice questions that include question descriptions, goals, options, and correct answers.
 They encapsulate valuable domain-specific knowledge that LLMs need to comprehend.
-The knowledge capacity of LLMs can be directly gauged by their performance on these test datasets.
-So far, static data-based benchmarks have successfully evaluated the relative knowledge capacity of LLMs in various domains [12].
-However, despite the high quality of expert-designed test data, static data-based benchmarks still face significant challenges in accurately quantifying the true knowledge capacity of LLMs due to their limited test scenarios and the unavoidable risk of data contamination.
-First, static data-based benchmarks depend on close-ended, single-formed questions for evaluation, which differs a lot to complicated real-world scenarios.
-For instance, instead of merely selecting an option, some users prefer asking LLMs to judge the correctness of options, while others may directly request the LLMs to generate an answer without providing an option list.
-Each possible prompt style may affect the performance of LLMs.
-Furthermore, Schaeffer et al.
-[13] have demonstrated that using nonlinear or discontinuous evaluation methods could lead to inaccurate evaluation results of LLMs’ capacity.
-Therefore, to genuinely evaluate the knowledge capacity of LLMs, it is necessary and feasible to test their ability to adjust to a range of complex scenarios that mirror real-world conditions.
-Second, static data-based benchmarks face an inherent risk of data contamination.
-These benchmarks publish test data online to ensure transparency and reproducibility in the evaluation process.
-However, this practice provides opportunities for LLM developers to maliciously fine-tune their models to "memorize" the test data, thereby undermining the benchmark’s integrity.
-This can result in an overestimation of the LLMs’ knowledge capacity and significantly compromise the reliability of the evaluation results.
-Efforts to prevent data contamination have primarily focused on detecting such contamination [14, 15] and generating new test data [16].
-Nonetheless, high-quality test data contains valuable knowledge that has not been fully leveraged.
-We contend that it is feasible to utilize this knowledge effectively to assess the true knowledge capacity of LLMs while mitigating the risk of data contamination.
-To achieve these goals, we propose PertEval, a knowledge-invariant perturbation-based evaluation toolkit for unveiling real knowledge capacity of LLMs in static data-based benchmarks.
-This idea stems from an analogy between LLM evaluation and humans’ educational assessment.
-Similar to LLM evaluation, close-ended questions are prevalent in human evaluation due to their versatility, cost effectiveness and precision of measurement [17], and also suffers from multiple challenges such as cheating [18] and a lack of variety [17].
-Solutions include but limited to item personalization [19] and the design of distractive options [20, 21].
-Inspired by this, PertEval includes human-like knowledge-invariant perturbations to transform test data to various complicated forms, and transition analysis methods to evaluate and trace the change of the performance of LLMs in different test scenarios.
-Knowledge-invariant perturbations consist of the rewriter LLM-based content-level perturbation for largely changing text features of questions to prevent data contamination, and a series of format-level perturbations for comprehensively covering possible test scenarios in real-world conditions.
-Transition analysis methods consist of not only evaluation metrics for measuring real knowledge capacity of LLMs, but also response pattern analysis for tracing the cause of the change of LLM’s performance, which probes in-depth weaknesses of the knowledge capacity of LLMs and provide insights for the refinement of LLMs.
-In experiments, we re-evaluate six state-of-the-art LLMs using PertEval.
-Transition analysis results first reveal a significant overestimation of the knowledge capacity of LLMs in MMLU, with an absolute 21% overestimation for GPT-4 and 35% overestimation for Gemini-1.0-Pro.
-Response pattern analyses further unveil that the overestimation for most LLMs is caused by a significant increment of the ratio of extra incorrect choices in response patterns, with an increment up to 9.7% for GPT-4 and 27.7% for Gemini-1.0-Pro.
-These results illustrate the potential rote memorization of the LLMs on correct options.
-Detailed transition analyses in terms of overall performance stability and correct response consistency reveal various weaknesses in the knowledge capacity of existing LLMs, providing guidance for their refinement.
-Given these insights, we posit that PertEval serves as an essential tool for evaluating the real knowledge capacity of LLMs when applying to any existing close-ended benchmarks, marking a significant step towards more trustworthy LLM evaluation.
+The knowledge capacity of LLMs can then be directly gauged by the performance on these test datasets.
+To date, static benchmarks have been essential in assessing the relative knowledge capacity of LLMs [12] because of their high quality and cost effectiveness.
+However, due to limited test scenarios [13] and the unavoidable risk of data contamination [14], these benchmarks still encounter significant challenges in accurately measuring the true knowledge capacity of LLMs.
+First, static benchmarks rely exclusively on close-ended questions in fixed formats for evaluation, which differs a lot from real-world scenarios [15].
+For instance, instead of merely selecting an option, some users prefer asking LLMs to judge the correctness of options, while others may directly request the LLMs to generate an answer without providing an option list [16].
+These varying prompting styles have an innegligible effect on performance [17].
+Furthermore, previous work [18] has demonstrated that nonlinear or discontinuous evaluation metrics could lead to inaccurate implications about LLM capacity.
+Therefore, to genuinely evaluate the knowledge capacity of LLMs, it is necessary to test their performance across a variety of scenarios that mirror real-world conditions.
+Second, in terms of data contamination [14], knowledge capacity benchmarks usually publish their test data online to ensure transparency and reproducibility (e.g., [6, 7]).
+However, this practice enables LLMs to memorize the test data during pre-training and alignment, which can lead to an overestimation of knowledge capacity, thus undermining the trustworthiness of evaluation results [19].
+Efforts to mitigate data contamination have primarily focused on detecting such contamination [14, 20] and generating new test data [21].
+Nonetheless, high-quality benchmark data contain rich knowledge that is valuable for evaluation.
+We suppose that it is feasible to utilize this knowledge effectively to assess the true knowledge capacity of LLMs while reducing the risk of data contamination.
+To achieve this goal, we introduce PertEval, an evaluation toolkit that utilizes knowledge-invariant perturbations on static benchmarks to unveil the real knowledge capacity of LLMs, as illustrated in Figure 1.
+This idea stems from an analogy with human educational assessment.
+Like LLM evaluation, close-ended questions are prevalent in human assessments due to their versatility, cost-effectiveness and precision of measurement [22].
+However, they also face challenges such as cheating [23] and a lack of variety [22].
+Solutions to these issues include item personalization [24] and the creation of distractive options [25, 26].
+Inspired by this, PertEval incorporates human-like knowledge-invariant perturbations to restate static test data to various forms and employs response consistency analyses to evaluate and trace the change of LLMs’ performance in different test scenarios.
+Specifically, knowledge-invariant perturbations consist of both the content-level perturbation that extensively writes questions to minimize data contamination and a series of format-level perturbations that comprehensively cover potential real-world test scenarios.
+Response consistency analyses encompass basic evaluation metrics to measure the real knowledge capacity of LLMs and response pattern analysis to investigate the causes behind changes in LLM performance.
+Consequently, PertEval could deeply probe the LLMs’ weaknesses in knowledge mastery and offer insights for their refinement.
+In experiments, we re-evaluate six representative LLMs using PertEval.
+Capacity metric results first reveal a significant overestimation of the knowledge capacity of LLMs by static benchmarks, with an absolute 25.8% overestimation for GPT-4 and 38.8% for Gemini-1.0-Pro on MMLU [6] (results of Gemini-1.0-Pro are available at Appendix D.2).
+Response pattern analyses further unveil that the performance decline with PertEval is mainly caused by an increase of selecting extra incorrect choices, with an increase of up to 9.7% and 27.7% for GPT-4 and Gemini-1.0-Pro, respectively.
+These findings highlight the potential for rote memorization of correct options by LLMs on static benchmarks.
+Detailed response consistency analyses in terms of overall performance stability and correct response consistency indicate various weaknesses in the knowledge capacity of existing LLMs, such as the vulnerability to content-level perturbation and the change of global text order.
+Overall, our contributions are as follows:
+  • We propose an evaluation toolkit that utilizes knowledge-invariant perturbations on close-ended evaluation benchmarks to unveil the real knowledge capacity of LLMs, a significant step towards more trustworthy LLM evaluation.
+  • We re-evaluate the knowledge capacity of six representative LLMs using PertEval.
+Evaluation results not only demonstrate significantly inflated performance of LLMs by static benchmarks, but also reveal LLMs’ uncertainty to specious knowledge and rote memorization to correct options.
+  • We demonstrate the vulnerability of various LLMs to different perturbation strategies in PertEval and provide insights for the refinement in terms of promoting LLMs’ knowledge capacity.
 2 Related Work
-Benchmarks for Knowledge Capacity Evaluation.
-Benchmarks for knowledge capacity evaluation could be classified into general knowledge benchmarks or professional knowledge benchmarks [12].
-General knowledge benchmarks aim to evaluate the general knowledge capacity of LLMs in all kinds of domains, such as MMLU [6], C-Eval[22] and ARC[8].
-On the other hand, professional knowledge benchmarks aim to deeply evaluate the professional capacity of LLMs in specific domains, such as MedMCQA [9] (Medicine), ScienceQA [23] (Science) and GSM-8K [11] (Mathematics).
-These benchmarks depend on professional multiple-choice questions to measure the capacity of LLMs.
-Trustworthiness of Knowledge Capacity Evaluation.
-Research about the trustworthiness of knowledge capacity evaluation are mainly from the test data acquisition perspective and the response generation perspective.
-For test data acquisition, the data contamination risk have been proved widely exist in test data [24, 25, 26].
-Recent works emphasize the detection of contamination [14, 15] and generating contamination-free test data [16, 27].
-For response generation, it has been observed that multiple text-related biases could affect LLMs’ performance [28, 29, 30, 31], such as option ordering bias [28] and selection bias [29].
-Debiasing methods like PriDe [14] serve as external plug-in for LLMs in multiple-choice question answering, which are inappropriate for evaluating real knowledge capacity.
-Adversarial Text Attack for LLMs.
-Adversarial text attacks aim to mislead language models to generate wrong outputs.
-Important research directions of them in the era of LLM include but not limited to jailbreak attack and text classification attack.
-Traditional text attacks could be classified into character-level [32, 33], word-level [34] and sentence level [35].
-Recently, LLM-based perturbations such as PromptAttack [36] have also been introduced.
-However, considering their cost and the comprehensiveness, existing text attack methods are insufficient for achieving our research goal.
+Knowledge Capacity Evaluation of LLMs Research works of the knowledge capacity evaluation LLMs consist of two lines, i.e., evaluation benchmarks and evaluation methodologies.
+The first line aim to design comprehensive and accurate benchmarks to quantize the knowledge capacity of LLMs.
+Evaluation benchmarks could be further classified into general or professional knowledge benchmarks [12].
+The first category aims to evaluate the general knowledge capacity of LLMs across a wide range of domains, such as MMLU [6], C-Eval [27] and ARC [8].
+The second category aims to deeply evaluate the professional capacity of LLMs in specific domains, such as MedMCQA [9] in medicine and ScienceQA [28] in science.
+These benchmarks depend on professional multiple-choice questions to quantitatively measure the capacity of LLMs.
+The other research line, i.e., evaluation methodologies, aim to improve the accuracy and truthfulness of evaluation via data-driven or model-driven techniques.
+Along this line, recent studies have emphasized data-contamination detection techniques [14, 20] and contamination-free test data generation [21, 29].
+In addition, psychometric-based techniques [30], dynamic evaluation techniques [31, 32], and perturbation-based evaluation techniques such as CheckList [33] and PolyJuice [34] have been proposed to comprehensively evaluate the capability of language models (LMs) from different aspects.
+However, these methodologies are unsuitable for knowledge capacity evaluation, and are hard to utilize valuable information in expert-designed datasets for trustworthy evaluation.
+Consistency of LMs The consistency of an LM denotes its behavior invariance under meaning-preserving alternations in its input [35], which is significant in natural language understanding.
+Since the emergence of pretrained language models (PLMs), massive efforts have been made to evaluate the consistency of LMs on various fields.
+Elazar et al.
+[35] revealed the poor consistency performance of PLMs on factual knowledge.
+Fierro and Søgaard [36] extended the study to multilingual PLMs and obtained similar findings.
+Jang et al.
+[37] proposed a taxonomy for various concepts of consistency and established a benchmark, BECEL, to evaluate the consistency of PLMs.
+As LLMs develop swiftly, Wang et al.
+[38] proposed to use self-consistency decoding to empower the Chain-of-Thought reasoning ability of LLMs.
+Recently, Rajan et al.
+[39] proposed KonTest, an autonomous evaluation framework that utilizes knowledge graphs to generate test samples, to probe the inconsistency in LLMs’ knowledge of the world.
+In summary, these research works view consistency as a research subject that needs to be measured or intervened.
+Differently, in our study, the consistency of LLMs is viewed as a technical method, i.e., a measurement for probing LLMs’ real knowledge capacity.
+Adversarial Text Attacks on LMs These techniques aim to mislead language models to yield wrong or toxic outputs with small text perturbations, which play an indispensable role in the research of robustness and safety of language models.
+Such efforts include but not limited to jailbreaking [40] and text classification attack [41].
+Text attacks could also be classified into character-level [42, 43], word-level [44] and sentence-level [45].
+Recently, LLM-based perturbations like PromptAttack [41] have also been introduced.
+However, considering their cost and the comprehensiveness, existing text attack methods are insufficient for the purpose of this work.
+Moreover, our PertEval is built upon a new concept of knowledge-invariant perturbation, which is an important supplement to this topic.
 3 Methodology
-Table 1: Standards of knowledge invariance
+3.1 Knowledge-invariant Perturbations
+We first present the two types of knowledge-invariant perturbations to restate static test questions. The knowledge-invariant property of these perturbations will be discussed in the next subsection.
+Content-level perturbation: knowledge-invariant paraphrasing.
+The goal of content-level perturbation is to substantially alter the phrasing of questions while retaining the original knowledge, thereby mitigating data contamination in the test data.
+The key challenge of such knowledge-invariant paraphrasing is to preserve the original knowledge while changing the statement as much as possible.
+A test question is a composite of sentences, with each provide either backgrounds, conditions or goals of the question.
+Therefore, to preserve knowledge of the original question, we propose a sentence-by-sentence paraphrasing algorithm using an LLM rewriter.
+Formally, let $q=(s_{1},s_{2},\ldots,s_{T})$ represent the question text, where $s_{t}$ denotes the $t$-th sentence (for $t=1,2,\ldots,T$).
+The semantic of $s_{t}$ (for $t\geq 2$) depends on its prerequisite sequence $(s_{1},\ldots,s_{t-1})$.
+To rewrite the entire question, an LLM rewriter is instructed to paraphrase sentence by sentence given each sentence and its original prerequisite sequences.
+Detailed descriptions of the paraphrasing algorithm, the prompt template for the rewriter LLM, and an example of the perturbation can be found in Appendix B.1.
+Format-level perturbation: question format refactoring.
+The objective of question format refactoring is to assess the robustness of LLMs’ knowledge capacity under complicated test conditions.
+To achieve this, we have developed a variety of format-level perturbation strategies to comprehensively evaluate the resilience of LLMs’ knowledge capacity in various test scenarios.
+  • Option permutation (OptionPerm).
+OptionPerm reorders the contents of options while maintaining the original order of option IDs.
+Its goal is to evaluate option ordering bias in the knowledge acquisition of LLMs.
+By default, OptionPerm reverses the order of option contents to completely disrupt their local sequence.
+  • Option format refactoring (OptionForm).
+OptionForm modifies the format of option IDs, such as by appending a right parenthesis to the end.
+This perturbation aims to assess the option format bias in the knowledge acquisition of LLMs.
+OptionForm may influence LLM performance by altering the dependency between different tokens within the options.
+  • Option ID shifting (OptionCaesar).
+OptionCaesar shifts the ASCII value of option IDs to change their character, which is similar to Caesar encryption.
+This technique aims to investigate the selection bias in the knowledge acquirement of LLMs, a phenomenon empirically observed in some models.
+By replacing common option IDs with less common ones (e.g., A/B/C/D $\to$ U/V/W/X), OptionCaesar allows us to observe whether the values of the IDs impact the LLMs’ performance.
+  • Question type changing (ChangeType).
+ChangeType converts a multi-choice question into a multi-judgment question.
+This perturbation aims to examine the question type bias in LLMs.
+Since the feasible solution space of a multi-choice question and the corresponding multi-judgement question is identical (given $N$ options, the size of the feasible solution space is $12^{N}-$ 1), an LLM that robustly acquire knowledge/skills in a question should be insensitive to ChangeType.
+  • Question position swapping (SwapPos).
+SwapPos switches the position of the question text and the options.
+It aims to evaluate the global ordering bias of LLMs.
+For rational human test-takers, SwapPos does not affect performance, as it does not alter the question content.
+However, SwapPos can significantly change the output distribution of self-regressive text generation models by disrupting the global ordering of input prompts.
+Examples of format-level perturbations are available at Appendix B.2.
+3.2 Knowledge Invariance Verification
+Table 1: Standards of knowledge-invariance scoring.
 Standard Name Standard Description
@@
 Statement Clarity The perturbed question should clearly present contexts, conditions and the target of the question.
-3.1 Knowledge-invariant Perturbation
-3.1.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
-The content-level perturbation aim to largely change the statement of questions while keeping the original knowledge to prevent data contamination of orginal test data.
-The key challenge of knowledge-invariant paraphrasing is how to preserve the original knowledge while change the statement as possible it can.
-To this end, we first propose standards of knowledge invariance, as presented in Table 1 for the checking of knowledge invariance.
-Next, we propose a sentence-by-sentence paraphrasing algorithm using an LLM rewriter.
-The motivation is that the logic of a question text is the composition of the logic of the sentence sequences.
-Formally, let $q=(s_{1},s_{2},\ldots,s_{T})$ be the question text, where $s_{t},t=1,2,\ldots,T$ denotes the $t$-th sentence.
-Then the semantic of $s_{t}(t\geq 2)$ depends on its prerequisite sequence $(s_{1},\ldots,s_{t-1})$.
-To rewrite the whole question text, an LLM rewriter is instructed to rewrite sentence by sentence given each sentence and its original context.
-The paraphrasing algorithm, the prompt template for the rewriter LLM and an example of the perturbation are available at Appendix C.1.
-3.1.2 Format-level Perturbation: Question Format Refactoring
-Question format refactoring aim to evaluate the robustness of LLMs’ knowledge capacity when confronting complicated test conditions.
-To this end, we design a variety of format-level perturbation strategies that cover most possible format bias of LLMs in real-world scenarios.
-Option Permutation (OptionPerm).
-OptionPerm reorders option contents while keeping the order of option IDs.
-It aims to evaluate the option ordering bias in the knowledge acquirement of LLMs.
-The default setting of OptionPerm is to reverse the order of option contents to completely destroy the local order of contents.
-When applying OptionPerm, the true answer of the question is also changed to be aligned with the perturbation result.
-Option Format Refactoring (OptionForm).
-OptionForm changes the format of option IDs (e.g., append a right parenthesis to the end of IDs).
-It aims to evaluate the option format bias in the knowledge acquirement of LLMs.
-OptionForm might affect the performance of LLMs by changing the dependency beteween different tokens in options.
-Option ID Shifting (OptionCaesar).
-OptionCaesar shifts the ASCII value of option IDs to change their character, which is similar to Caesar encryption.
-It aims to evaluate the selection bias in the knowledge acquirement of LLMs, whose existence has been empirically proved in some LLMs.
-By applying OptionCaesar, we replace common option IDs (e.g., A/B/C/D) by uncommon option IDs so that we can observe whether the value of IDs affect the performance of LLMs.
-Question Type Changing (ChangeType).
-ChangeType changes a multiple choice question to a multiple judgement question.
-It aims to evaluate the question type bias of LLMs.
-Since the feasible solution space of a multiple choice question and the corresponding multiple judgement question is identical (given $N$ options, the size of the feasible solution space is $(2^{N}-1)$), an LLM that robustly acquire knowledge/skills in a question should be insensitive to ChangeType.
-Therefore, it is a significant perturbation for LLMs’ knowledge acquirement evaluation.
-Question Position Swapping (SwapPos) SwapPos swaps the position of question texts with options.
-It aims to evaluate the global ordering bias of LLMs.
-For rational human test-takers, SwapPos does not change their performance on questions because merely move question text to after the options does not change semantic information and knowledge of the question.
-However, although token-level knowledge is also kept in SwapPos, this perturbation still might be a difficult challenge for LLMs, because it destroys the global ordering of input prompts, which might largely change the distribution of output of self-regression text generation models.
-3.2 Knowledge Invariance Checking
-LLM-based Knowledge Invariance Scoring.
-This method checks the knowledge invariance of perturbations from the perspective of LLM’s perception.
-Specifically, we first choose a superior LLM (e.g., gpt-4-turbo) as the referee.
-Then, give the test dataset $D_{dual}=\{(q_{i},q_{i}^{\prime})|i=1,2,\ldots,|D|\}$, we construct the scoring prompt based on the prompt template, standards of knowledge invariant perturbations (see Table 1) and the criteria of scores for knowledge invariance judgement (see Table 8 in Appendix C.1) for each pair of original and perturbed questions.
-After collecting the output score of the referee LLM from the scoring prompt, the average score is calculated as the final result.
+We next investigate the knowledge-invariance property of the proposed perturbations.
+Knowledge invariance scoring.
+This approach checks the knowledge invariance of perturbations from the perspective of humans’ and LLMs’ perception.
+Specifically, based on previous works in the measurement of semantic similarity [46] and characteristics of close-ended benchmarks, we first propose standards of knowledge-invariance scoring, as shown in Table 1.
+Next, we recruit professional human volunteers and utilize superior LLMs, such as claude-3-sonnet, respectively, to serve as the referee to rate knowledge invariance scores.
+Given a set of original and perturbed question pairs $D_{dual}=\{(q_{i},q_{i}^{\prime})\mid i=1,2,\ldots,|D|\}$, we construct scoring prompts based on a predefined template, knowledge invariant standards, and the scoring criteria for knowledge invariance judgement (see Table 9 in Appendix B.1).
+After collecting the output scores from the referee LLM using these scoring prompts, we calculate the average score as the final result.
 Testing on mastered questions for LLMs.
-Different from the scoring method that is input-oriented, this method evaluates the output-oriented knowledge invariance.
-The motivation is that if a perturbation is knowledge-invariant, an LLM’s performance on questions whose knowledge it really mastered should be the same on the original and the perturbed version.
-The challenge is that mastered questions for LLMs are hard to find.
-To this end, we propose a substitute solution that utilizes questions that most LLMs can correctly answer to represent the set of mastered questions for an LLM.
-Then an LLM test-taker (gpt-4-turbo in experiments) is required to answer the original and perturbed versions of these questions to validate the knowledge invariance of the perturbation.
-In a word, the test-based knowledge invariance checking procedure can be summarized as follows:
+This approach evaluates knowledge invariance based on the output performance of LLMs.
+The rationale here is that if a perturbation is knowledge-invariant, an LLM’s performance on questions that it has truly mastered should remain consistent between the original and perturbed versions.
+Identifying mastered questions for LLMs, however, poses a challenge.
+To this end, we propose using questions that most LLMs can correctly answer as a proxy for mastered questions.
+An LLM, such as gpt-4-turbo, is then required to answer both the original and perturbed versions of these questions to validate the knowledge invariance of the perturbation.
+In summary, the test-based knowledge invariance checking procedure involves the following steps:
   1. 1.
-    Given a set of LLM test-takers and the evaluation dataset $D$, collect question samples that all LLMs answer correctly to construct the mastered question set $D_{simple}$.
+    Given a set of LLM test-takers and the evaluation dataset $D$, construct the mastered question set $D_{simple}\subset D$ consisting of questions that all LLMs could answer correctly.
   2. 2.
-    Apply a perturbation to each of sample of $D_{simple}$ to obtain the perturbed dataset $D_{simple}^{\prime}$.
+    Apply a perturbation to each sample in $D_{simple}$ to create the perturbed dataset $D_{simple}^{\prime}$.
   3. 3.
-    For each LLM, evaluate its performance on $D_{simple}^{\prime}$.
-If there exists at least one LLM whose performance on $D_{simple}^{\prime}$ has no significant difference with that on $D_{simple}$, then the perturbation is knowledge-invariant.
-This method focuses more on the influence of perturbation on the text generation process of LLMs.
-Compared to knowledge invariance scoring operations, this methods is more suitable for evaluating the knowledge invariance of content-level perturbations.
-3.3 Transition Analysis For Measuring Real Knowledge Capacity
-The transition analysis aims to measure the real knowledge capacity of LLMs from their responses to benchmark datasets.
-In this part, we first propose the metric of real knowledge capacity.
-Next, we propose two aspects for detailed analyses of LLMs’ weaknesses in knowledge capacity.
+    Assess the performance of a knowledgeable LLM on both $D_{simple}$ and $D_{simple}^{\prime}$.
+If the test results show no significant difference in performance, the perturbation can be considered knowledge-invariant.
+This method emphasizes the impact of perturbations on the question-answer process of LLMs.
+In comparison to the scoring test, this approach is more appropriate for evaluating the knowledge invariance of content-level perturbations.
+We present detailed results of both tests in Experiments.
+3.3 Response Consistency Analyses for Measuring Real Knowledge Capacity
+We further devise a suite of response consistency analyses that, by comparing performance on raw vs.
+perturbed test sets, unveils the real knowledge capacity of LLMs.
+These include a metric for quantifying calibrated knowledge capacity and the more fine-grained ones helpful for revealing the vulnerability of LLMs to different perturbation strategies.
 Metric of real knowledge capacity.
 To measure real knowledge capacity of LLMs, we propose the Consistent Accuracy (ACC@Consist) as the evaluation metric.
-The motivation is that if a question and its corresponding knowledge is truly mastered by an LLM, then the LLM’s performance should be consistent on any perturbed version of the question, including the original version and the most complicated version.
-To this end, let $M(\cdot)$ denotes the response function of an LLM.
-Let $x=(q_{x},y_{x})\in X$ denotes a test question, where $q_{x}$ denotes the question text and $y_{x}$ denotes the correct answer(s).
-Let $\sigma^{*}:X\to X$ denotes the most complicated composite knowledge-invariant perturbation.
-Let $D=\{x_{1},x_{2},\ldots,x_{n}\}$ denotes the test set.
+The rationale is that if an LLM has truly mastered a question and its underlying knowledge, its performance should remain consistent across all versions of the question, including both the original and the most complex perturbed versions.
+Formally, let $M(\cdot)$ denotes the response function of an LLM.
+Let $x=(q_{x},y_{x})\in X$ represent a test question, where $q_{x}$ and $y_{x}$ denote the question text and the correct answer(s) of $x$.
+Further let $\sigma^{*}:X\to X$ be the most complicated composite knowledge-invariant perturbation and $D=\{x_{1},x_{2},\ldots,x_{|D|}\}$ the raw test set.
 Then ACC@Consist is defined as:
 […]
-Here $I(\cdot)$ denotes the indicator function.
-Aspect 1: Overall Performance Stability.
-In terms of overall performance, we expect LLMs that have robustly acquired knowledge/skills required by benchmark datasets should have a stable performance when confronting knowledge-invariant perturbations.
-To this end, we propose Performance Drop Rate (PDR) to measure the over performance stability of LLMs given knowledge-invariant perturbations:
-$\text{PDR}(M,D,\sigma)=\frac{1}{|D|}\sum_{x\in D}I[M(q_{\sigma(x)})=y_{\sigma(x)}]-I[M(q_{x})=y_{x}].$ (2)
-Here $M$ denotes the LLM, $D$ and $D^{\prime}$ denotes the original and the perturbed dataset respectively.
-$PDR$ essentially equals to the discrepancy between the LLM’s accuracy on $D$ and that on $D^{\prime}$.
-If $PDR<0$, then the perturbation decreases the overall performance of LLMs.
-Then the LLM does not robustly acquire knowledge/skills.
-To further obtain reliable conclusions, we further make Wilcoxon signed-rank test for original and perturbed question sample pairs.
-Table 2: Performance Transition Matrix
-Number Correct after perturbation $\sigma$ Incorrect after perturbation $\sigma$
-Correct before perturbation $\sigma$ $CC$ $IC$
-Incorrect before perturbation $\sigma$ $IW$ $CW$
-Aspect 2: Correct Response Consistency.
-To this end, we propose the Recall of Performance (ROP) metric.
-Its name originates from an analogy to the classical Recall score.
-To define ROP, we first propose the performance transition matrix, which is similar to the confusion matrix, as shown in Table 2.
-Here the first capital character of each variable (i.e., C/I) denotes “consistent” or “inconsistent”.
-The second capital character of each variable (i.e., C/N) denotes “correct” or “wrong”.
-For example, the variable $CC$ means the number of “consistent correct” before and after the perturbation.
-Next, similar to recall score, ROP is defined as the ratio of consistent correct in the number of all correct samples before perturbation, i.e.,
-$\text{ROP}(M,D,\sigma)=\frac{CC}{CC+IC}.$ (3)
-Tha range of ROP is in $[0,1]$.
-The higher the ROP, the better the correct response consistency given the perturbation.
-It should be noticed that, unlike PDR, there does not exist a “threshold” for ROP to measure the correct response consistency.
-Rather, the ROP serves as a score of correct response consistency, which can be utilized to compare the performance of different LLMs and strategies.
-(a) Knowledge invariance score rated by GPT-4.
-(b) Logarithm of Levenshtein distance.
-Figure 1: LLM-based knowledge invariance scoring using systematic sampling (interval = 10).
+Here $I(\cdot)$ is the binary indicator function. Taking a step further, we implement a response pattern analysis that uncovers the causes of the discrepancy between ACC@Consist and the original ACC.
+Overall performance stability.
+In terms of performance stability, we expect that LLMs which have robustly acquired the knowledge and skills required by benchmarks should exhibit stable performance when faced with knowledge-invariant perturbations.
+To evaluate this, we propose Performance Drop Rate (PDR) as a metric to measure the overall performance stability of LLMs under knowledge-invariant perturbations.:
+$\text{PDR}(M,D,\sigma)=\frac{1}{|D|}\sum_{x\in D}\Big{(}I[M(q_{\sigma(x)})=y_{\sigma(x)}]-I[M(q_{x})=y_{x}]\Big{)}.$ (2)
+Here $D$ denotes the original test dataset and $\sigma$ is a perturbation strategy.
+Essentially, $PDR$ measures the discrepancy between the LLM’s accuracy on $D$ and the perturbed dataset.
+When $PDR<0$, the perturbation decreases the overall performance of an LLM, indicating that it does not robustly acquire knowledge and skills.
+To obtain more reliable conclusions, we further conduct Wilcoxon signed-rank test for original and perturbed question sample pairs.
+Correct response consistency.
+We also adopt the Recall of Performance (ROP) metric to measure correct response consistency.
+The term “ROP” draws an analogy to the classical recall score.
+Let $CC$ denote the number of responses consistently correct before and after perturbation, and $IC$ the number of responses initially correct and later incorrect after perturbation (see Appendix B.3 for detailed demonstration).
+Similar to the recall score, ROP is defined as the ratio of consistent correct responses to the total number of correct responses before perturbation, i.e.,
+$\text{ROP}(M,D,\sigma)=CC/(CC+IC)$ (3)
+The value of ROP is within $[0,1]$, with higher ROP indicating better correct response consistency under perturbation $\sigma$.
+It should be noticed that, unlike PDR, there does not exist an absolute threshold for ROP to evaluate goodness.
+Rather, ROP serves as an index of correct response consistency, which can be utilized to compare the performance of different LLMs under perturbations.
+Table 2: Knowledge invariance $\uparrow$ rated by human scorers. Four independent scores from different human scorer groups are presented in ascending order for each cell.
+Method C-Math W-History P-Psychology P-Medicine
+PromptAttack 2.3/2.4/3.0/3.9 2.2/2.2/2.3/2.8 1.3/2.4/2.8/2.8 1.6/2.5/3.5/3.6
+PertEval (ours) 3.6/3.8/3.9/3.9 3.7/4.1/4.1/4.3 4.3/4.4/4.5/4.7 4.2/4.3/4.4/4.6
+Table 3: Knowledge invariance $\uparrow$ rated by superior LLMs.
+Values (a/b/c) in each cell denotes the average knowledge invariance score rated by gpt-4-turbo, claude-3.5-sonnet, and llama-3.1-405b, respectively.
+Method C-Math W-History P-Psychology P-Medicine
+PromptAttack 3.2/3.6/3.6 3.2/3.3/3.7 3.9/3.9/3.7 4.1/4.3/4.2
+PertEval (ours) 3.8/3.9/4.0 4.0/4.2/4.0 4.0/4.4/4.0 4.1/4.4/4.0
+Figure 2: (left) Perturbation-wise knowledge invariance $\uparrow$ by gpt-4-turbo (systematic sampling, interval = 10); (right) edit distances between original question and perturbed questions.
 Table 4: Results of testing on mastered questions - Performance Drop Rate (PDR) of overall performance stability testing on mastered questions using gpt-4-turbo.
@@
 SwapPos $0$0.1149 -0.0636 -0.0488 -0.0568
-4 Experiment
+4 Experiments
 Datasets.
-In the experiment, to comprehensively cover evaluation domains and subjects while keeping the evaluation cost, we select test data in College Mathematics (C-Math), World History (W-History), Professional Psychology (P-Psychology) and Professional Medicine (P-Medicine) from the Massive Multitask Language Understanding (MMLU) benchmark, with each representing a supercategory in STEM, Humanities, Social Sciences and Other respectively.
-Statistics of selected datasets are presented in Table 9 in Appendix D.
+We choose the test data in College Mathematics (C-Math), World History (W-History), Professional Psychology (P-Psychology) and Professional Medicine (P-Medicine) from the Massive Multitask Language Understanding (MMLU) [6] for evaluation.
+This is a trade-off between comprehensively covering evaluation domains/subjects and keeping the evaluation cost affordable.
+Note that each dataset represents a supercategory in STEM, Humanities, Social Sciences, and Other fields, respectively.
+The statistics of the selected datasets are detailed in Table 11 in Appendix C.1.
 Large Language Models.
-Based on existing LLM leaderboards like OpenCampass [^1] https://rank.opencompass.org.cn/ and considering both the popularity and timeliness of LLMs, we select six representative LLMs for evaluation.
-These LLMs include close-sourced (gpt-4-turbo, gpt-3.5-turbo, gemini-1.0-pro, glm-3-turbo) and open-sourced (mistral-7b-instruct-v0.2, llama-3-8b-instruct).
-4.1 Checking Knowledge Invariance for Perturbations
-Knowledge invariance checking results are presented in Figure 1 and Table 3.
-For LLM-based knowledge invariance scoring (Figure 1), we compare the score of proposed perturbations with the baseline PromptAttack[36] and find the excellent average score ($>4.5$) for each proposed perturbation.
-We also calculate the Levenshtein distance (edit distance) of perturbed and original questions for each perturbation, and we find that the traditional edit distance is not suitable for measuring knowledge invariance because of a lack of the correlation between edit distance and knowledge invariance score.
-For testing on mastered questions (Table 3), we find that most perturbations obtain an average PDR closed to zero, which indicates their consistent performance on the perturbed and the original data.
-The PDR of SwapPos on World History and Professional Psychology is less than -0.05 although it does not change any knowledge-relevant information based on our design.
-A possible explanation is that SwapPos changes the global ordering of question prompts, thus affects the text generation of these self-regression-based LLMs.
-More detailed result analyses are available at Appendix D.1
-4.2 Real Knowledge Capacity Evaluation (Transition Analysis Part 1)
-Figure 2: Real knowledge capacities measured by ACC@Consist with composite knowledge-invariant perturbation.
-Metrics are averaged on datasets.
-ACC@Original denotes accuracy from the original data.
-ACC@Perturb denotes accuracy from the perturbed data.
-ACC@Consist denotes ratio of consistent correct responses in both the original and the perturbed data.
-In PertEval, each knowledge-invariant perturbation strategy focuses on a specific question transformation.
-To construct the most challenging test scenario for LLMs and obtain reliable and conprehensive evaluation results, we compose all knowledge-invariant perturbations together to obtain the composite perturbation to quantify LLMs’ real knowledge capacity.
-Then the real knowledge capacity of LLMs is measured by ACC@Consist.
-Evaluation results are shown in Figure 2.
-Based on these results, we obtain several findings in terms of LLMs’ knowledge capacity evaluation:
+Based on leaderboards like OpenCompass [47] and considering both the popularity and timeliness of LLMs, we select six representative LLMs for evaluation.
+These LLMs include close-sourced (gpt-4-turbo [48], gpt-3.5-turbo [49], gemini-1.0-pro [50] and glm-3-turbo [51]) and open-sourced (mistral-7b-instruct-v0.2 [52] and llama-3-8b-instruct [53]) models.
+4.1 Knowledge Invariance Verification for Perturbations
+We first investigate the knowledge invariance property of our perturbations using the two methods developed in Section 3.2.
+Employing the LLM-based knowledge invariance scoring method, we compare the scores of the proposed perturbations with those by PromptAttack [41] as a baseline.
+Experiment Setup.
+For knowledge invariance scoring, 10 samples at equal intervals for each subject are used for scoring.
+Since there are four subjects and two perturbation strategies (1 baseline + 1 PertEval) for each question, we have 10*4*2 = 80 question pairs in total.
+To ensure that each question pair have four independent scores, eight human scorers are engaged in human-based scoring.
+Human scorers are equally divided into four groups.
+Each group independently scores for all 80 question pairs.
+Therefore, each human scorer within a group scores for 40 shuffled samples.
+For LLM-based scoring, we choose gpt-4-turbo, claude-3.5-sonnet, and llama-3.1-405b as scorers.
+For fine-grained perturbation-wise scoring, we choose gpt-4-turbo as the scorer.
+Result Analysis.
+Overall knowledge invariance scores are presented in Table 2 and 3.
+PertEval outperforms the baseline, and the score mostly exceeds 4.0, the borderline of knowledge-invariance.
+Scores of both PromptAttack and PertEval on C-Math do not exceed 4.0.
+This is because many samples in College Mathematics of MMLU are short mathematical reasoning questions that have many mathematical symbols and statements.
+Indeed, in many STEM subjects, requirements for knowledge capacity and reasoning ability often mix together in a test question.
+This points out considerable potential for future development in robust LLM evaluation in STEM subjects.
+Next, the results presented in Figure 2 show that the average scores of our perturbations approach 5.0 (perfectly knowledge-invariant) in most cases, consistently outperforming those from PromptAttack.
+We also present the Levenshtein distance (i.e., edit distance) between the perturbed and original questions under each perturbation.
+It turns out that the traditional edit distance is not suitable for measuring knowledge invariance due to its lack of correlation with knowledge invariance.
+Within mastered question testing (results depicted in Table 4), we observe that most perturbations achieve an average PDR closed to zero, indicating consistent performance on the perturbed and the original data.
+The PDR of SwapPos on World History and Professional Psychology is less than -0.05 despite that it does not alter any knowledge-relevant information by design.
+A possible explanation is that SwapPos changes the global ordering of question prompts, thus affecting the token generation of these self-regressive LLMs.
+More detailed analyses of results could be found in Appendix D.
+[…]
+4.2 Real Knowledge Capacity Evaluation
+Each knowledge-invariant perturbation in PertEval targets a specific question restatement strategy.
+To construct the most challenging test scenario and obtain reliable and comprehensive evaluation results, we compose all perturbations to create a composite perturbation.
+The real knowledge capacity of LLMs is then quantified by ACC@Consist before and after the composite perturbation.
+The results, shown in Figure 3, yield several insights regarding the evaluation of LLMs’ knowledge capacity.
 1.
-The knowledge capacity of LLMs is significantly overvalued in static dataset-based benchmarks.
-Comparing the ACC@Original with ACC@Perturb and ACC@Consist for each LLM, we observe a significant drop of LLMs’ performance on perturbed dataset.
-Especially for gpt-4-turbo, gpt-3.5-turbo and gemini-1.0-pro, the change exceeds 20%.
-Even the most powerful selected LLM, gpt-4-turbo, gains only ACC@Consist as 0.5551.
-These results indicate a large gap between evaluated knowledge capacity of LLMs in static data-based benchmarks and real knowledge capacity of LLMs in complicated scenarios.
+Overvaluation on static benchmarks.
+The knowledge capacity of LLMs is significantly overvalued on static benchmarks.
+By comparing ACC@Original with ACC@Perturb and ACC@Consist for each LLM, we observe a notable performance drop when using the perturbed dataset.
+For instance, the performance of gpt-4-turbo, gpt-3.5-turbo, and gemini-1.0-pro decreases by more than 25%.
+Even the most capable LLM in our selection, gpt-4-turbo, achieves an ACC@Consist of only 0.555.
+These findings reveal a substantial gap between the evaluated knowledge capacity of LLMs on static datasets and their real knowledge capacity in complex scenarios.
 2.
-Every selected LLM has its own consistently mastered knowledge.
-This result is deduced from a comparison between LLMs’ performance and pure guess.
-Given $k$ options and a single answer, the expected ACC@Consist is $1/k^{2}$.
-Therefore, the expected of ACC@Consist for pure guess is $0.06251/4^{2}=$ 0.0625 in the experiment ($k=4$)222see Appendix D.2 for detailed demonstration..
-We observe from Figure 2 that ACC@Consist values of all selected LLMs are significantly larger than 0.0625, which means that each selected LLM indeed has its own consistently mastered knowledge, although most of them master less than a half (ACC@Consist < 0.5) of the total knowledge.
-4.3 Response Pattern Analysis (Transiton Analysis Part 2)
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 3: Response patterns of gpt-4-turbo in all datasets.
-We conduct response pattern analysis to explore how perturbations affect the performance of LLMs, and we discover that knowledge-invariant perturbations affect LLMs by increasing the ratio of extra multiple choices.
-As presented in Figure 3, the main reason why the performance of gpt-4-turbo drops dramatically is that they select extra incorrect options more frequently in the perturbed data (gpt-4-turbo: $12\%\to 21.7\%$).
-These LLMs tend to select extra incorrect options besides selecting correct options.
-A possible explanation of this phenomenon is that the LLMs memorize correct options, while lack the ability to recognize and filter out incorrect options, which means that these LLMs do not really master knowledge of the questions.
-Therefore, besides the ability to distinguish correct answers, we should also emphasize enhancing large language models’ ability to recognize and filter out incorrect answers when facing options in real-world problems.
-More detailed response pattern analysis results are available at Appendix D.3.
-4.4 Overall Performance Stability (Transition Analysis Part 3)
-This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of overall performance.
-Evaluation results of overall performance stability are shown in Table 4.
-Here each value in the table’ cell is the macro PDR, i.e., the average value of PDRs separately calculated on each dataset.
-The Wicoxon signed-rank test is applied to the micro PDR, i.e., the overall PDR calculated on the all-in-one dataset.
-Specifically, let $s=I(M(q_{x})=y_{x})\in\{0,1\}$ and $s^{\prime}=I(M(q_{\sigma(x)})=y_{\sigma(x)})\in\{0,1\}$ respectively denotes model $M$’s score on the original and the perturbed questions, the alternative hypothesis of the Wilcoxon signed-rank test is $H_{a}:s>s^{\prime}$, or, $H_{a}:s^{\prime}-s<0$.
-From Table 4, we can both analyze for each column or for each row.
-For columns (strategy-aspect), we first observe that all selected LLMs shows negative macro PDRs given the content-level perturbation, the knowledge-invariant paraphrasing (KnInvPara).
-In addition, for close-sourced LLMs, the micro PDR is significantly negative, which means that the these LLMs significantly lack the content-level knowledge acquirements robustness in these datasets.
-Another observation is that all the LLMs are highly sensitive to the SwapPos perturbation.
-We suppose the reason is that, unlike other format-level perturbations that mostly focus on local format, SwapPos changes the global format of questions, which highly destroys self regression-based LLMs’ text generation process, even though this perturbation is entirely knowledge-invariant.
+Robustly mastered knowledge.
+Each LLM does have robustly mastered certain knowledge.
+This conclusion is drawn by comparing the LLMs’ performance to pure guessing.
+With $k$ options and a single correst answer, the expected ACC@Consist for random guessing is $1/k^{2}$, which equals to 0.0625 for $k=4$ in our experiments (see Appendix D.1 for detailed demonstration).
+We observe from Figure 3 that the ACC@Consist values of all tested LLMs are significantly higher than 0.0625, indicating that each LLM has indeed mastered some knowledge consistently.
+However, most LLMs have mastered less than half of the total knowledge, as indicated by ACC@Consist values below 0.5.
+Figure 4: Response patterns of gpt-4-turbo. Left: original data; Right: perturbed data.
+4.3 Response Pattern Analysis
+We further conduct a response pattern analysis to examine how perturbations affect the performance of LLMs.
+We find that knowledge-invariant perturbations affect LLMs by increasing the ratio of selecting extra incorrect options.
+As shown in Figure 4, the main reason for the significant performance drop in gpt-4-turbo is their increased frequency of selecting additional incorrect options on the perturbed data (i.e., $12\%\to 21.7\%$).
+These LLMs tend to select extra incorrect options alongside the correct ones.
+One possible explanation for this phenomenon is that the LLMs indeed lack the ability to distinguish uncertain knowledge and filter out distracting options, while they correctly answer these questions in the original dataset via rote memorization.
+In other words, these LLMs do not truly master the knowledge behind the questions.
+Therefore, it is crucial to not only enhance LLMs’ ability to identify correct answers but also bolster their capability to recognize and eliminate incorrect answers when confronted with distracting choices in real-world scenarios.
+More detailed response pattern analysis results can be found in Appendix D.2.
+4.4 Overall Performance Stability
+Table 5 presents the overall performance stability results, in which each number is the macro-PDR, i.e., the average of PDRs separately calculated on each dataset.
+We also apply the Wicoxon signed-rank test to the micro-PDR, which is the PDR calculated across all datasets combined.
+Specifically, let $s=I(M(q_{x})=y_{x})\in\{0,1\}$ and $s^{\prime}=I(M(q_{\sigma(x)})=y_{\sigma(x)})\in\{0,1\}$ denote model $M$’s score on the original and perturbed questions, respectively.
+The alternative hypothesis of the Wilcoxon signed-rank test is $H_{a}:s>s^{\prime}$, or equivalently, $H_{a}:s^{\prime}-s<0$.
+We can analyze the results of Table 5 from both a column-wise and row-wise perspective.
+From the column-wise perspective, we first observe that all selected LLMs have negative macro-PDRs given the content-level perturbation, i.e., knowledge-invariant paraphrasing (KnInvPara).
+Additionally, the micro-PDR of close-sourced LLMs is significantly negative, meaning that these LLMs lack robustness in content-level knowledge acquisition for these datasets.
+Another prominent observation is that all LLMs are highly sensitive to the SwapPos perturbation.
+We hypothesize that this sensitivity is due to the global format change introduced by SwapPos, which disrupts the text generation process of self-regressive LLMs, even though the perturbation is entirely knowledge-invariant.
+From the row-wise perspective, we find that our perturbations cause performance drop in most cases, e.g., at least 0.04 and up to 0.1 on average, indicating the universal vulnerability of current generation LLMs.
 Table 5: Macro PDR $\uparrow$ and hypothesis test results of Micro PDR of LLMs w.r.t. perturbation.
 […]
-gpt-4-turbo -$0∗$0208∗∗ -0.0136 -0.0294∗∗ -0.0210 -0.1117∗∗ -0.0468
-gpt-3.5-turbo -0.0275∗∗ -0.0042 -0.1767∗∗ -0.0396∗∗ -0.1736∗∗ -0.1943∗∗ -0.1047
+gpt-4-turbo -$0∗$0208∗∗ -0.0136 -0.0294∗∗ -0.0210 -0.1117∗∗ -0.0438
+gpt-3.5-turbo -0.0275∗∗ -0.0042 -0.1767∗∗ -0.0396∗∗ -0.1736∗∗ -0.1943∗∗ -0.1027
 gemini-1.0-pro -0.0558∗∗ +0.0121 +0.0125 +0.0030 -0.1310∗∗ -0.1532∗∗ -0.0521
-glm-3-turbo -$0∗$0190 -0.1397∗∗ -0.0118 +0.0522 -0.2142∗∗ -0.0667
-mistral-7b-instruct-v0.2 -0.0264 -$0$0.2789∗∗ +0.0793 -0.0844∗∗ -0.1275∗∗ -0.0763
-llama-3-8b-instruct -0.0336∗∗ -0.0091 -0.0939∗∗ -0.0368∗∗ -0.2920∗∗ -0.1814∗∗ -0.1074
-AVG -0.0456 -0.0087 -0.0994 0.0003 -0.0596 -0.1335
+glm-3-turbo -$0∗$0190 -0.1397∗∗ -0.0118 +0.0522 -0.2142∗∗ -0.0616
+mistral-7b-v0.2 -0.0264 -$0$0.2789∗∗ +0.0793 -0.0844∗∗ -0.1275∗∗ -0.0763
+llama-3-8b -0.0336∗∗ -0.0091 -0.0939∗∗ -0.0368∗∗ -0.2920∗∗ -0.1814∗∗ -0.1078
+AVG -0.0411 -0.0102 -0.1151 -0.0059 -0.1083 -0.1637
   • ∗∗: The Micro PDR is significantly negative in the Wilcoxon signed-rank test ($\alpha=0.01$).
@@
 […]
-gpt-4-turbo 0.8798 0.9063 0.9601 0.9349 0.9221 0.7995 0.8949
-gpt-3.5-turbo 0.8230 0.7728 0.5602 0.7801 0.5194 0.5610 0.6631
+gpt-4-turbo 0.8798 0.9063 0.9601 0.9349 0.9221 0.7995 0.9005
+gpt-3.5-turbo 0.8230 0.7728 0.5602 0.7801 0.5194 0.5610 0.6694
 gemini-1.0-pro 0.7968 0.7995 0.9553 0.9226 0.6788 0.6362 0.7982
-glm-3-turbo 0.7164 0.6756 0.5943 0.7884 0.6816 0.4026 0.6288
-mistral-7b-instruct-v0.2 0.6543 0.6202 0.1339 0.7679 0.5128 0.4060 0.5158
-llama-3-8b-instruct 0.8055 0.7433 0.7541 0.8102 0.3971 0.5209 0.6676
-AVG 0.7522 0.7624 0.6673 0.8323 0.6858 0.6009
-4.5 Correct Response Consistency (Transition Analysis Part 4)
-This analysis aim to unveil the vulnerability of LLMs to different perturbations in terms of the consistency of correct responses.
-Experimental results of ROP are presented in Table 5.
-First, in terms of LLMs, the correct response consistency of different LLMs vary a lot.
-Among them, gpt-4-turbo performs best with the average macro ROP as 0.8949, which is a big gap compared to other LLMs.
-We also observe that gemini-1.0-pro performs well especially on the OptionForm and OptionCaesar strategies (Macro PDR > 0.9), which can detect the selection bias of LLMs.
-Without prerequisite information, we suppose that special effort have been made in the training of gemini-1.0-pro to overcome selection bias.
-For other LLMs, their correct response consistencies are vulnerable to almost all knowledge-invariant perturbations, which means that their correct answer in the original datasets are less reliable.
-On the other hand, the effect of different strategies on correct response consistency also vary a lot.
-Similar to results in overall performance stability, SwapPos is the most powerful strategy that influences LLMs’ correct response consistency, which exposes the inherent flaw of LLMs in defending global order-level perturbation.
-5 Conclusion
-In this paper, we proposed PertEval, an evaluation toolkit for measuring large language models’ real knowledge capacity using human-like knowledge-invariant perturbations and transition analysis methods.
-Based on the well-defined knowledge perturbation strategies and metrics in transition analysis, we revealed the overestimated knowledge capacity of six state-of-the-art LLMs on the MMLU benchmark.
-Response pattern analyses discovered that PertEval retains LLMs’ uncertainty to specious knowledge and revealed their rote memorization to correct options.
-Further transition analysis by PertEval illuminated the vulnerability of different LLMs to different perturbations and could guid the development of refinement.
-Finally, we discussed ways to enhance the knowledge capacity of LLMs using PertEval.
+glm-3-turbo 0.7164 0.6756 0.5943 0.7884 0.6816 0.4026 0.6432
+mistral-7b-v0.2 0.6543 0.6202 0.1339 0.7679 0.5128 0.4060 0.5159
+llama-3-8b 0.8055 0.7433 0.7541 0.8102 0.3971 0.5209 0.6719
+AVG 0.7793 0.7530 0.6597 0.8340 0.6186 0.5544
+4.5 Correct Response Consistency
+The results of correct response consistency using ROP are presented in Table 6.
+There is significant variation in ROP across different LLMs.
+Among them, gpt-4-turbo performs the best, with an average macro-ROP of 0.9, which is a substantial advantage over other LLMs.
+Gemini-1.0-pro also performs well, particularly with the OptionForm and OptionCaesar strategies (both having a Macro-ROP > 0.9).
+Other LLMs, however, are vulnerable to almost all knowledge-invariant perturbations, which means that their performance on the original datasets are less reliable.
+On the other hand, the effect of different strategies on correct response consistency varies considerably.
+Similar to the results for performance stability, SwapPos is the most influential strategy affecting LLMs’ correct response consistency, which exposes the inherent flaw of LLMs in defending global order perturbation.
+5 Discussion
+Conclusion.
+The trustworthiness of capacity evaluation is, and will always be one of the most significant issues in the development of LLMs.
+In pursuit of this goal, we proposed an evaluation toolkit, PertEval, for unveiling LLMs’ real knowledge capacity on close-ended benchmarks.
+We discovered not only a substantial overestimation of the knowledge capacity of LLMs by static benchmarks, but also LLMs’ uncertainty to distinguish specious knowledge without rote memorization.
+Detailed response consistency analyses revealed the vulnerability of LLMs to various knowledge-invariant perturbations, especially the content-level KnInvPara and the format-level SwapPos, which could inspire further refinement for LLMs’ knowledge mastery.
+Indeed, the ultimate goal of this study is not only trustworthy evaluation, but to propel the development of LLMs’ knowledge capacity.
+Our primary experiments have verified the feasibility of using PertEval-generated data to improve LLM robustness to various test conditions via supervised fine-tuning (see Appendix D.3).
+We believe that PertEval marks a significant step toward more trustworthy LLM evaluation.
+Limitations & Future Work.
+The first limitation of this work is its scope regarding suitable benchmarks.
+PertEval is suitable for close-ended benchmarks and cannot be directly applied to open-ended ones.
+The challenges stem from the variety of open-ended questions and the difficulty of their objective evaluation.
+Future research could explore how the concept of knowledge-invariant perturbation can be adapted to improve the trustworthiness of open-ended evaluations.
+The second limitation of PertEval is the lack of controllability, a common issue in LLM-based text generation.
+Due to the complexity of LLMs, it is challenging to precisely control specific characteristics in paraphrased texts while keeping knowledge-invariance.
+In the future, we aim to achieve more fine-grained controllability in PertEval, such as quantitatively adjusting paraphrasing tone and length, to obtain more continuous and nuanced evaluation results of LLMs.
+Besides, all conclusions of this study are based on experiments conducted on a subset of MMLU due to budget constraints.
+We are seeking economic methods to expand our experiments and further substantiate our findings.
+Finally, the risk of adaptive optimization against PertEval exists as it is an open-sourced toolkit.
+However, the diversity of perturbations and the cost of adaptive optimization could effectively reduce the influence of adaptive optimization.
+By constructing a PertEval open evaluation platform, we can regularly develop and change perturbation strategies to prevent this procedure.
+Acknowledgements
+This research was supported by grants from the Joint Research Project of the Science and Technology Innovation Community in Yangtze River Delta (No.
+2023CSJZN0200), the National Natural Science Foundation of China (No.
+62337001), the Fundamental Research Funds for the Central Universities, and the Alibaba Research Intern Program.
+We thank all the volunteers for their massive efforts in supporting our experiments, including Yi Cheng, Xiaowen Zhang, and Anyu Chen at Alibaba Cloud Computing, Dingchu Zhang at Nanjing University, Yuyang Xu at Zhejiang University, and Yangyang Wang at Shandong University.
 Appendix
 Appendix A Related Work
-A.1 LLMs’ Knowledge Capacity Evaluation Benchmarks
+This section complements related works especially in LLM evaluation for Section 2.
+A.1 Benchmarks for Knowledge Capacity Evaluation
 LLMs’ knowledge capacity evaluation plays a fundamental yet significant role in LLM evaluation tasks.
@@
 Howver, how to avoid such bias to evaluate LLMs’ real knowledge capacity remains a research question.
-Appendix B Discussion about PertEval
-The ultimate goal of PertEval is not only to unveil the real knowledge capacity of LLMs, but also to guide the refinement of LLMs in enhancing knowledge capacity.
-To this end, we discuss potential methods and future directions for enhancing LLMs’ knowledge capacity with PertEval:
-Supervised fine-tuning with perturbed data.
-PertEval can be utilized to generate various forms of multiple-choice question data.
-For knowledge that requires memorization such as history and laws, new data generated by PertEval can be used to enhance the ability of LLMs to adjust to various complicated test scenarios.
-Indeed, we conduct a primary experiment using llama-3-8b-instruct (see Appendix D.4) and reveal the feasibility of this solution.
-Extending PertEval to open-ended test data.
-The current PertEval adapts only to multiple-choice questions considering their precision of measurement and cost effectiveness.
-To further evaluate the knowledge capacity of LLMs in all kinds of domains, a future direction is to extend PertEval to open-ended data types such as open question answering.
-Potential challenges in this direction include but not limited to how to define the correctness of output answer given perturbations and how to ensure knowledge invariance in complicated questions such as solving math application questions.
-Appendix C Details of Methodology
-C.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
+Appendix B Details of Methodology
+B.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
 The algorithm of knowledge-invariant paraphrasing using LLM rewriter is presented in Algorithm 1.
@@
 You should only output the rewrited sentence without any extra content. Expected similarity score: {similarity_score} Context: {context} Sentence: {sentence} Your output:
-Figure 4: Prompt template for the rewriter LLM.
+Figure 5: Prompt template for the rewriter LLM. The expected similarity score is used to control parphrasing of the rewriter LLM, which is set to 0.6 in all experiments.
 Table 7: An example of knowledge-invariant paraphrasing of a test question. Texts surrounded by angular brackets are invisible in question prompts input to the LLM test-taker.
@@
 <# Goal> Then, the linear transformation $T$ will map the point (2, 1) to
-C.2 Format-level Perturbation: Question Format Refactoring
+B.2 Format-level Perturbation: Question Format Refactoring
 Examples of question format refactoring are shown in Table 8.
@@
 Figure 7: Prompt template for the LLM-based knowledge invariance scoring (Part 2/2).
-Appendix D Details of Experiments
+B.3 Correct Response Consistency
+Similar to the confusion matrix, terms in ROP can be written in a response consistency matrix, as shown in Table 10.
+Here “$C$”/“$I$” in the first position denotes “consistently”/“inconsistently”.
+“$C$”/“$W$” in the second position denotes “correct”/“wrong”.
+For instance, $CC$ denotes consistently correct.
+Table 10: Response transition matrix.
+Number Correct after perturbation $\sigma$ Wrong after perturbation $\sigma$
+Correct before perturbation $\sigma$ $CC$ $IC$
+Wrong before perturbation $\sigma$ $IW$ $CW$
+Appendix C Experiment Setup
+C.1 Details of Dataset
 Statistics of datasets are presented in Table 11.
@@
 # Tokens per P 129.52$\pm$29.19 392.27$\pm$127.62 125.11$\pm$31.63 233.58$\pm$66.60
+C.2 Evaluation Setting
+Model Setting.
+For close-sourced LLMs (gpt-4-turbo, gpt-3.5-turbo, glm-3-turbo, gemini-1.0-pro), we use their APIs for evaluation.
+The temperature parameter of each model is set to 0.2 to reduce output uncertainty.
+For open-sourced LLMs (llama-3-8b-instruct, mistral-7b-instruct-v0.2), we locally deploy their huggingface version to our own server and utilize default settings for evaluation.
+Environment Setting.
+All experiments are completed on a Linux server with Intel(R) Xeon(R) Platinum 8269CY CPUs @ 2.50GHz and one NVIDIA A100 GPU (40G).
+GPUs are used for deploying and fine-tuning open-sourced models.
+The version of Python is 3.9.
+The version of the torch package is 2.2.0.
+The version of the transformers package is 4.39.3.
+Prompt Setting.
+In our experiments, we use prompt templates to generate prompts of questions for evaluation.
+Examples using prompt templates are shown in Figure 8 and 9.
+After getting model outputs, we transform them to json data for further analyses and utilize regular expressions to enhance the robustness of transformation.
+Details are available at our code.
+Supervised Fine-tuning Setting.
+In Appendix D.3, we fine-tune llama-3-8b-instruct using PertEval-generated data to explore the utility of PertEval in enhancing the knowledge capacity of LLMs.
+The fine-tuning process is done using LoRA [65] on single GPU implemented in LLaMA-Factory [66].
+The batch size is 1, since the number of training data is 237 (High School World History) + 272 (Professional Medicine) = 479.
+The learning rate is 5e-5.
+The number of training epoch is 3.
+Please select the correct option(s) from the following options given the question: Question: Let k be the number of real solutions of the equation $e^{x}+x-2=0$ in the interval [0, 1], and let n be the number of real solutions that are not in [0, 1].
+Which of the following is true?
+Options: A k = 0 and n = 1 B k = 1 and n = 0 C k = n = 1 D k > 1 Your output must strictly follow this format: {"answer": <the list of selected options, e.g., ["A", "B", "C", "D"]>} Your output:
+Figure 8: Prompt template for multiple-choice question answering.
+Please judge whether each of the options is correct or incorrect given the question: Question: Let k be the number of real solutions of the equation $e^{x}+x-2=0$ in the interval [0, 1], and let n be the number of real solutions that are not in [0, 1].
+Which of the following is true?
+Options: A k = 0 and n = 1 B k = 1 and n = 0 C k = n = 1 D k > 1 Your output must strictly follow this format: {"A": <"True" or "False">, "B": <"True" or "False">, "C": <"True" or "False">, "D": <"True" or "False">} Your output:
+Figure 9: Prompt template for multiple-judgement question answering (ChangeType).
 Figure 10: Accuracy score of LLMs in original datasets.
 The gpt-4-turbo@Mastered denotes the accuracy of gpt-4-turbo on the mastered question set for checking knowledge invariance, which is defined in Section 3.2.
-D.1 Checking the Knowledge Invariance of Perturbation Strategies
-In this part, we utilize the proposed knowledge invariance checking methods to measure the validity of perturbations in terms of knowledge invariance.
-The prompt template for the LLM-based knowledge invariance scoring is presented in Figure 5 and 6.
+C.3 Experiment Cost
+Costs in # Tokens. Overall, we conduct the following experiment steps that require token costs:
+  1. E0.
+    Generating PertEval benchmark.
+  2. E1.
+    Knowledge-invariance validation - obtain knowledge invariance score.
+  3. E2.
+    Main experiment - obtain ACC@Consist.
+  4. E3.
+    Fine-grained perturbation analysis - obtain ROP and PDR.
+Exact costs w.r.t. experiments are presented as follows:
+Table 12: Exact token costs w.r.t. experiments.
+Index # Input tokens # Output tokens
+E0 3,542,896 374,967
+E1 776,353 101,675
+E2 3,248,124 190,476
+E3 9,343,200 336,996
+Costs in USD.
+For proprietary LLMs, the cost is from the pricing strategy of LLM APIs (proportional to # input/output tokens).
+For open-sourced LLMs, the cost originates from the pricing strategy of servers for LLM deployment (proportional to the running time of servers).
+We count overall costs in USD per experiment, as shown in Table 13.
+Table 13: Costs in USD per experiment.
+Exp Index Price (USD)
+E0 41.79
+E1 10.81
+E2 13.09
+E3 51.47
+Total 117.16
+Appendix D Verifying the Knowledge Invariance of Perturbation Strategies
+In this part, we utilize the proposed knowledge invariance verification methods to measure the validity of perturbations in terms of knowledge invariance.
+The prompt template for the LLM-based knowledge invariance scoring is presented in Figure 6 and 7.
 Results of LLM-based knowledge invariance scoring are shown in Figure 2.
@@
 We observe from Table 4 that SwapPos can decrease the performance of gpt-4-turbo even in mastered questions although we have demonstrated the knowledge invariance of the perturbation in Figure 2, Comparing PDRs on different datasets and statistics of different datasets, we observe that it is highly correlated with the number of tokens per question.
-D.2 Probing LLMs’ Consistent Knowledge Capacity
+D.1 Probing LLMs’ Consistent Knowledge Capacity
 Proposition D.1.
@@
 […]
-$E[\text{ACC@Consist}]=E[X/n]=\frac{1}{n}E[X]=1/k^{2}$ (4)
+$E[\text{ACC@Consist}]=E[Z/n]=\frac{1}{n}E[Z]=1/k^{2}$ (4)
 Then the proof is completed. ∎
-D.3 Response Pattern Analysis for LLMs in PertEval
+D.2 Response Pattern Analysis for LLMs in PertEval
 In this part, we analyze response pattern for LLMs in the original and perturbed data, which can provide us a deep insight into the influence of knowledge-invariant perturbations on LLMs’ question answering procedure.
-Specifically, response patterns consist of correct choice, invalid choice, extra multiple choice, wrong single choice and wrong multiple choice.
-Their definition are listed in the following:
-  • Correct choice: The LLM selects and only selects all correct options. The ratio of correct choices is equvalent to micro accuracy.
-  • Invalid choice: The LLM selects none of options. This could be caused by the LLM’s wrong reasoning process or invalid output format.
-  • Extra multiple choice: The LLM selects not only all correct options, but also extra incorrect options. This means that the LLM fails to filter out incorrect options from the option list.
-  • Wrong single choice: The LLM selects only a single option, and the option is wrong. This means that the LLM fails to recognize the correct option in the option list.
-  • Wrong multiple choice: The LLM selects multiple options which do not cover all correct options.
-Given the response pattern taxonomy, we counted the ratio of response patterns in original and perturbed data for all selected LLMs, as shown in Figure 3, 8, 9, 10, 11 and 12.
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 8: Response patterns of gpt-3.5-turbo in all datasets.
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 9: Response patterns of glm-3-turbo in all datasets.
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 10: Response patterns of gemini-1.0-pro in all datasets.
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 11: Response patterns of llama-3-8b-instruct in all datasets.
-(a) Original data
-(b) Perturbed data with composite perturbation
-Figure 12: Response patterns of mistral-7b-instruct-v0.2 in all datasets.
-D.4 How to Enhance LLMs’ Knowledge Capacity Using PertEval?
+The definition of response patterns are listed in the following:
+  • Single Correct Selection: The LLM selects and only selects the correct option. The ratio of correct choices is equvalent to micro accuracy.
+  • Single Wrong Selection: The LLM selects only a single option, and the option is wrong. This means that the LLM fails to recognize the correct option in the option list.
+  • Invalid Selection: The LLM selects none of options. This could be caused by the LLM’s wrong reasoning process or invalid output format.
+  • Multiple Selection Including the Correct One: The LLM selects not only the correct option, but also extra incorrect options.
+This means that the LLM fails to filter out incorrect options from the option list.
+  • Multiple Selection w/o the Correct One: The LLM selects multiple options which do not cover the correct option.
+Given the response pattern taxonomy, we counted the ratio of response patterns in original and perturbed data for all selected LLMs, as shown in Figure 4, 11, 12, 13, 14 and 15.
+Figure 11: Response patterns of gpt-3.5-turbo in all datasets. Left: original data. Right: perturbed data.
+Figure 12: Response patterns of glm-3-turbo in all datasets. Left: original data. Right: perturbed data.
+Figure 13: Response patterns of gemini-1.0-pro in all datasets. Left: original data. Right: perturbed data.
+Figure 14: Response patterns of llama-3-8b-instruct in all datasets. Left: original data. Right: perturbed data.
+Figure 15: Response patterns of mistral-7b-instruct-v0.2 in all datasets. Left: original data. Right: perturbed data.
+D.3 How to Enhance LLMs’ Knowledge Capacity Using PertEval?
 The goal of PertEval is not only to evaluate large language models’ consistent knowledge capacity, but to help LLMs enhance their knowledge capacity in real-world applications.
@@
 Considering the numerous number of knowledge domains and the issue of cost and efficiency, we suppose training knowledge expert models for specific knowledge domains is a good choice for striking a balance between knowledge capacity and training efficiency.
-Table 10: PDR of LLaMA-8B-Instruct before and after finetuning on perturbed datasets.
-Underlined datasets are used for finetuning.
+Table 14: PDR of LLaMA-8B-Instruct before and after fine-tuning on perturbed datasets.
+Underlined datasets are used for fine-tuning.
 CT denotes ChangeType.
@@
 SP denotes SwapPos.
-For example, F(CT) means finetuning LLaMA-8B-Instruct on world history and professional medicine datasets perturbed by ChangeType.
+For example, F(CT) means fine-tuning LLaMA-8B-Instruct on world history and professional medicine datasets perturbed by ChangeType.
 […]
-ChangeType Original -0.2300 -0.3418 -0.2467 -0.3493 -0.2920 -0.1998
+ChangeType None -0.2300 -0.3418 -0.2467 -0.3493 -0.2920 -0.1998
 F(CT) -0.0700 +0.0759 +0.0196 +0.0257 +0.0128 -0.0868
 F(CT+KP) -0.0500 +0.0422 +0.0082 +0.0074 +0.0020 +0.0019
-SwapPos Original -$0$0.2110 -0.1944 -0.2500 -0.1814 -0.2867
+SwapPos None -$0$0.2110 -0.1944 -0.2500 -0.1814 -0.2867
 F(SP) +0.0100 -0.0675 -0.1095 -0.0882 -0.0638 +0.0246
 F(SP+KP) -$0$0.1350 -0.1029 -0.1176 -0.0964 -0.1065
-KnInvPara Original +0.0200 -0.0802 -0.0163 -0.0478 -0.0311 -0.0328
+KnInvPara None +0.0200 -0.0802 -0.0163 -0.0478 -0.0311 -0.0328
 F(KP) -$0$0.0253 -0.0212 $0$0.0216 -0.0188
@@
 F(SP+KP) -$0$0.0675 -0.0212 -0.0184 -0.0343 -0.0303
+Checklist
+  1. 1.
+    For all authors…
+    1. (a)
+      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
+    2. (b)
+      Did you describe the limitations of your work? [Yes] See Section 5.
+    3. (c)
+      Did you discuss any potential negative societal impacts of your work?
+[No] Our work mainly focuses on techniques for LLM evaluation.
+We proposed a novel evaluation toolkit for enhancing existing evaluation benchmark rather proprosing than a new dataset.
+Therefore, there is no risk for this work to have negative societal impacts caused by online data collection, such as the invasion of users’ privacy.
+    4. (d)
+      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
+  2. 2.
+    If you are including theoretical results…
+    1. (a)
+      Did you state the full set of assumptions of all theoretical results? [Yes]
+    2. (b)
+      Did you include complete proofs of all theoretical results? [Yes] Due to page limitation, proofs are available at Appendix D.1.
+  3. 3.
+    If you ran experiments (e.g. for benchmarks)…
+    1. (a)
+      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
+[Yes] The URL to our code and data is available at abstract.
+    2. (b)
+      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Due to page limitation, details are available at Appendix C.
+    3. (c)
+      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)?
+[No] It is too costly and timely expensive to conduct experiments for multiple times given the huge number of settings (assuming there are 6 models, 6 perturbations, 4 datasets and 10 running times, there are at least 6*6*4*10 = 1,440 groups of experiments).
+Substitutionally, in experiments, we set the temperature parameter to all LLMs as a relative small value ($\tau=0.2$) to reduce the uncertainty of LLMs’ outputs.
+We also open-sourced our code to ensure reproducibility.
+    4. (d)
+      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)?
+[Yes] A description of computational resources is available at Appendix C.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
+    1. (a)
+      If your work uses existing assets, did you cite the creators? [Yes] We used LLMs (both APIs of close-sourced LLMs and weights of open-sourced LLMs) and open-sourced datasets in our work.
+    2. (b)
+      Did you mention the license of the assets? [No]
+    3. (c)
+      Did you include any new assets either in the supplemental material or as a URL? [Yes] The URL of the code and data of our work is available at abstract.
+    4. (d)
+      Did you discuss whether and how consent was obtained from people whose data you’re using/curating?
+[N/A] We did not collect private data from people.
+Instead, all APIs, model weights and evaluation data used in this work, are published online and accessible to everyone.
+    5. (e)
+      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content?
+[N/A] Data we used in our work is about knowledge capacity evaluation, and does not contain personally identifiable information.
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects…
+    1. (a)
+      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
+    2. (b)
+      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
+    3. (c)
+      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]